### v2

# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

Abstract.
We describe the rings of invariants for the finite orthogonal groups of plus type in odd characteristic acting on the defining representations. We also describe the invariants of the corresponding Sylow subgroups in the defining characteristic. In both cases we construct minimal algebra generating sets and describe the relations among the generators. Both rings of invariants are shown to be complete intersections and thus are Cohen-Macaulay. We expect the techniques we use will generalise to give a systematic computation for rings of invariants for all of the finite classical groups in odd characteristic.
Key words and phrases:
modular invariant theory, finite classical groups, orthogonal group
1991 Mathematics Subject Classification:
Primary 13A50; Secondary 20F55
1. Introduction
The fundamental problem in the invariant theory of finite groups is to determine the ring of invariants of a representation of a finite group. The celebrated theorem of Shepherd–Todd and Chevalley proves that such a ring of invariants in characteristic zero is a polynomial ring if and only if the group action is generated by pseudo-reflections (group elements which fix pointwise a hyperplane). An excellent survey of the characteristic zero results and theory in characteristic zero is the article by Stanley ([22]).
In positive characteristic the situation is much more complex. In 1911, L.E. Dickson [12] gave an explicit description of the ring of invariants of the general linear group over any finite field. In the past few decades there has been significant progress in determining the rings of invariants of the defining representations of the finite classical groups. The rings of invariants for the symplectic groups were computed by David Carlisle and Peter Kropholler in the 1990s, see [1, §8.3]. The invariants in the unitary case were computed by Huah Chu and Shin-Yao Jow [9]. For the orthogonal groups, while there are computations in a number of special cases (see [7], [8], [10], [17] and [21] ) there is no general result. While almost all of the defining representations for these groups are generated by pseudo-reflections, the rings of invariants are rarely polynomial rings.
We work over the finite field $\mathbb{F}_{q}$ of order $q$ and characteristic $p$ where $p$ is an odd prime. We compute the ring of invariants for the finite orthogonal group ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ of plus type acting on its defining representation $V$ of dimension $n=2m$.
As an intermediate result, we compute the ring of invariants for a Sylow $p$-subgroup. For both of these rings we construct a minimal set of algebra generators and describe generators for the ideal of relations. Both of these rings are complete intersections, see Theorems 4.6 and 5.9, and are thus Cohen-Macaulay. We note that for modular representations, rings of invariants are rarely Cohen-Macaulay, let alone complete intersections.
For a Sylow $p$-subgroup $P$ of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ we find a set $\mathcal{H}_{0}$ consisting of $n$ orbit products of linear forms such that $\mathcal{H}_{0}$ forms a homogeneous system of parameters. Using a family of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ invariants $\xi_{0},\xi_{1},\dots,\xi_{n-2}$, we show that
$\mathbb{F}_{q}[V]^{P}=\bigoplus_{\gamma\mid\Gamma_{0}}\mathbb{F}_{q}[\mathcal{H}_{0}]\gamma$
where the sum is over all monomials $\gamma$ dividing
$\Gamma_{0}=\prod_{i=0}^{n-3}\xi_{i}^{e_{i}}\text{ where }e_{i}=q^{m-1-\lfloor i/2\rfloor}-1\ .$
In particular, $\mathbb{F}_{q}[V]^{P}$is Cohen-Macaulay and this implies that $\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ is Cohen-Macaulay. We also give a Khovanskii basis for the Sylow invariants (see the definition at the end of the next section).
We show that there are homogeneous orthogonal invariants $d_{1,m},d_{2,m},\ldots,d_{m,m}$ such that =$\mathcal{H}=\{\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{m,m}\}$ is a homogeneous system of parameters. Furthermore, we show that
$\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}=\bigoplus_{\gamma\mid\Gamma}\mathbb{F}_{q}[\mathcal{H}]\gamma$
where the sum is over all monomials $\gamma$ dividing
$\Gamma=\prod_{i=m}^{n-2}\xi_{i}^{q^{n-i-1}-1}\ .$
For the orthogonal invariants, the ring is generated by two elements as an algebra over the Steenrod algebra (see Section 3.2 and Theorem 8.1).
We approximate $\mathbb{F}_{q}[V]^{P}$ using the lead term algebra since there exists a finite Kovanskii basis. However we do not expect this to happen for the invariants of the full orthogonal group. Thus we approximate $\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ by the associated graded algebra arising from a filtration given by a certain valuation.
For both the orthogonal group and its Sylow $p$-subgroup, we express the ring of invariants as a free module over a homogeneous system of parameters with basis given by all monomial factors of a single monomial. Such a basis is called a block basis (see [4]). Since the rings are Cohen-Macaulay, having block bases implies that these rings are both complete intersections.
We believe that the methods we use here generalise to give a systematic computation for rings of invariants for all of the finite classical groups, at least for odd characteristic; this is work in progress. For background on the finite classical groups we suggest [23]. For background on invariant theory see [1], [5], [11] or [20].
2. The Setting
For a symmetric bilinear form $\beta$ on a vector space $V$, the orthogonal group ${\rm O}(\beta)$ is the subgroup of the the general linear group ${\rm GL}(V)$ consisting of linear transformations $g$ satisfying $\beta(gv,gu)=\beta(v,u)$ for all $u,v\in V$. In this paper we work over the finite field $\mathbb{F}_{q}$ of order $q$ and characteristic $p$ where $p$ is an odd prime.
In this context, if the dimension of $V$ is even, there are two equivalence classes of symmetric non-degenerate bilinear forms, one of plus type and one of minus type (see Chapter 11 of [23]). In this paper we focus on the plus type. We expect the generalisation of our methods to the minus type to be relatively straightforward.
We denote the dimension of $V$ by $n=2m$. The quadratic form $\xi_{0}$ associated to the bilinear form $\beta$ is defined by $\xi_{0}(v):=\beta(v,v)/2$ for $v\in V$. We choose an ordered basis for $V$, $[e_{1},\ldots,e_{m},f_{m},\ldots,f_{1}]$, so that $\beta(e_{i},f_{i})=1$, $\beta(e_{i},e_{j})=\beta(f_{i},f_{j})=0$ and, if $i\not=j$, $\beta(e_{i},f_{j})=0$, in other words, the basis consists of $m$ orthogonal hyperbolic pairs (see pages 56 and 139 of [23]). We note that $m$ is the Witt index and the basis vectors are isotropic. We let $[y_{1},\ldots,y_{m},x_{m},\ldots,x_{1}]$ denote the dual basis and observe that the quadratic form is given by
$\xi_{0}=y_{1}x_{1}+\cdots+y_{m}x_{m}.$
The right action of ${\rm GL}(V)$ on $V$ induces a left action on the dual $V^{*}$ given by $(\phi\cdot g)(v)=\phi(g\cdot v)$ for $\phi\in V^{*}$, $g\in{\rm GL}(V)$ and $v\in V$. This action extends to an action by algebra automorphisms on $\mathbb{F}_{q}[V]$, the symmetric algebra on $V^{*}$. Choosing a basis for $V$ allows us to identify ${\rm GL}(V)$ with the matrix group ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$ and identify $\mathbb{F}_{q}[V]$ with the polynomial algebra on $n=2m$ variables
$S_{m}:=\mathbb{F}_{q}[y_{1},\ldots,y_{m},x_{m},\ldots,x_{1}].$
[…]
${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}=\{g\in{{\text{GL}_{{n}}(\mathbb{F}_{q})}}\mid\xi_{0}\cdot g=\xi_{0}\}.$
Its order is given by
$|{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}|=2q^{m(m-1)}(q^{m}-1)\prod_{i=1}^{m-1}(q^{2i}-1),$
see [23, page 141]. For notational simplicity, we write $G_{m}:={{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. The ring of invariants is
$S_{m}^{G_{m}}=\{f\in S_{m}\mid f\cdot g=f,\,\forall g\in G_{m}\}.$
Clearly $\xi_{0}\in S_{m}^{G_{m}}$. For $i>0$, define $\xi_{i}:=y_{1}^{q^{i}}x_{1}+y_{1}x_{1}^{q^{i}}+\cdots+y_{m}^{q^{i}}x_{m}+y_{m}x_{m}^{q^{i}}$. Since taking the $q^{th}$ power is $\mathbb{F}_{q}$-linear, $\xi_{i}\in S_{m}^{G_{m}}$. To avoid cluttering the notation we suppress the $m$ from the notation for the $\xi_{i}$.
For an integral domain $A$, let $\mathcal{F}(A)$ denote the field of fractions. The following is due to Carlisle and Kropholler [6].
Theorem 2.1.
$\mathcal{F}(S_{m}^{G_{m}})=\mathbb{F}_{q}(\xi_{0},\ldots,\xi_{n-1})$.
Throughout, we order the variables taking $y_{1}>y_{2}>\cdots>y_{m}>x_{m}>\ldots>x_{1}$. There are many ways of extending this ordering to a monomial order on $S_{m}$. We will use both the lexicographic and the graded reverse lexicographic (referred to as grevlex) orders. For $f\in S_{m}$, we denote the lead term of $f$, with respect to whichever order is being considered, by $\operatorname{LT}(f)$. For a subalgebra $A\subset S_{m}$, $\operatorname{LT}(A)$ denotes the algebra generated by the lead terms of the elements of $A$. A Khovanskii basis for $A$ is a subset of $A$, say $\mathcal{B}$, such that $\{\operatorname{LT}(f)\mid f\in\mathcal{B}\}$ is a generating set for $\operatorname{LT}(A)$. A Khovanskii basis is a nice generating set for $A$. Khovanksii bases were previously known as SAGBI bases. For more on Khovanskii(SAGBI) bases see [5, §5.1].
3. Missing Invariants
Define $\overline{V}:=V\otimes\overline{\mathbb{F}}_{q}$ where $\overline{\mathbb{F}}_{q}$ is the algebraic closure of $\mathbb{F}_{q}$. For $X\subset S_{m}$, the variety associated to $X$ is given by
$\mathcal{V}(X):=\{v\in\overline{V}\mid f(v)=0,\,\forall f\in X\}\subseteq\overline{V}.$
Let $(S_{m}^{G_{m}})_{+}$ denote the augmentation ideal, i.e., the span of the homogeneous invariants of positive degree. Since $G_{m}$ is a finite group, $\mathcal{V}((S_{m}^{G_{m}})_{+})=\{\underline{0}\}$. Recall that $[e_{1},\ldots,e_{m},f_{m},\ldots,f_{1}]$ is the ordered basis for $V$ dual to the ordered basis for $V^{*}$ given by $[y_{1},\ldots,y_{m},x_{m},\ldots,x_{1}]$. In the following, we write $\mathcal{V}(\xi_{0},\ldots,\xi_{i})$ for $\mathcal{V}(\{\xi_{0},\ldots,\xi_{i}\})$.
Theorem 3.1.
For $s\geq m-1$,
$\mathcal{V}(\xi_{0},\ldots,\xi_{s})=\bigcup_{g\in G_{m}}g\cdot\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m}\}.$
Proof.
Observe that the variety $\mathcal{V}(\xi_{0},\ldots,\xi_{s})$ is closed under scalar multiplication since it is defined by homogeneous polynomials. Let $W_{m}$ denote $\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m}\}$. Since $\xi_{i}(v)=0$ for $v\in W_{m}$ and $\xi_{i}\in S_{m}^{G_{m}}$,
$\bigcup_{g\in G_{m}}g\cdot W_{m}\subseteq\mathcal{V}(\xi_{0},\ldots,\xi_{s}).$
Therefore, it is sufficient to prove that for $v\in\mathcal{V}(\xi_{0},\ldots,\xi_{m-1})$, there exists $g\in G_{m}$ such that $gv\in W_{m}$.
The proof is by induction on $m$. For $m=1$, we have $x_{1}(v)y_{1}(v)=0$. If $x_{1}(v)\not=0$, then take $g$ to be the transposition which interchanges $x_{1}$ and $y_{1}$.
For $m>1$, define $\bar{v}:=v-y_{m}(v)e_{m}-x_{m}(v)f_{m}$, let $\bar{\xi_{i}}$ denote the restriction of $\xi_{i}$ to $\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m-1},f_{m-1},\ldots f_{1}\}$ and identify $G_{m-1}$ with the point-wise stabiliser of ${\rm Span}_{\mathbb{F}_{q}}\{e_{m},f_{m}\}$ in $G_{m}$.
If $x_{m}(v)=0$ then $\bar{\xi_{i}}(\bar{v})=0$ for $i=0,\ldots,m-1$. By induction, there is an element $g\in G_{m-1}<G_{m}$ with $g\bar{v}\in W_{m-1}$. Observe that $gv\in W_{m}$.
For $u\in\overline{V}$ we use $\operatorname{Fr}^{i}(u)$ to denote the $i^{th}$ iteration of the Frobenius map on $u$. Recall that $\beta(u,u)=2\xi_{0}(u)$ and note that, for $i>0$, $\beta(u,\operatorname{Fr}^{i}(u))=\xi_{i}(u)$.
If $x_{m}(v)\not=0$ then we scale $v$ so that $x_{m}(v)=1$ and define $w:=v-\operatorname{Fr}(v)$. Observe that
$\xi_{0}(w)$ $=$ $\beta(v-\operatorname{Fr}(v),v-\operatorname{Fr}(v))/2=\left(\beta(v,v)+\beta(\operatorname{Fr}(v),\operatorname{Fr}(v))-2\beta(v,\operatorname{Fr}(v))\right)/2$
$=$ $\xi_{0}(v)+\xi_{0}(v)^{q}-\xi_{1}(v)=0$
and, for $i>0$,
$\xi_{i}(w)$ $=$ $\beta(v-\operatorname{Fr}(v),\operatorname{Fr}^{i}(v-\operatorname{Fr}(v)))=\beta(v-\operatorname{Fr}(v),\operatorname{Fr}^{i}(v)-\operatorname{Fr}^{i+1}(v))$
$=$ $\beta(v,\operatorname{Fr}^{i}(v))-\beta(\operatorname{Fr}(v),\operatorname{Fr}^{i}(v))-\beta(v,\operatorname{Fr}^{i+1}(v))+\beta(\operatorname{Fr}(v),\operatorname{Fr}^{i+1}(v))$
$=$ $\xi_{i}(v)-\xi_{i-1}(v)^{q}-\xi_{i+1}(v)+\xi_{i}(v)^{q}=0.$
Therefore $\xi_{i}(w)=0$ for $i=0,\ldots,m-2$. Since $x_{m}(w)=0$, we have $\bar{\xi_{i}}(\bar{w})=0$ for $i=0,\ldots,m-2$ and so by induction there is an element $g\in G_{m-1}<G_{m}$ with $g\bar{w}\in W_{m-1}$ and $gw\in W_{m}$. Hence $(x_{j}-x_{j}^{q})(gv)=x_{j}(gw)=0$ for $j=1,\ldots,m$ and $x_{j}(gv)\in\mathbb{F}_{q}$. Since $g$ stabilises ${\rm Span}_{\mathbb{F}_{q}}\{e_{m},f_{m}\}$ , we have $x_{m}(gv)=1$. For convenience, define $c_{j}=x_{j}(gv)\in\mathbb{F}_{q}$ for $j=1,\ldots,m-1$ and let $h$ denote the linear transformation given by $y_{j}h=y_{j}$ for $j=1,\ldots,m-1$, and $y_{m}h=y_{m}+\sum_{j=1}^{m-1}c_{j}y_{j}$, and $x_{j}h=x_{j}-c_{j}x_{m}$ for $j=1,\ldots,m-1$ and $x_{m}h=x_{m}$. Observe that $h\in G_{m}$ and, since $\xi_{0}(gv)=0$, we have $y_{m}(hgv)=y_{m}(gv)+\sum_{j=1}^{m-1}c_{j}y_{j}(gv)=0$. Let $\sigma$ denote the transposition which exchanges $x_{m}$ and $y_{m}$. We then have $x_{m}(\sigma hgv)=0$ and we can apply the induction argument as above. ∎
It follows from this theorem that we are missing at least $m$ invariants.
3.2. Steenrod Operations
The Steenrod operations are cohomology operations widely used in algebraic topology. They can be sensibly defined on $S_{m}$ and restrict to operations on $S_{m}^{G_{m}}$. Thus applying a Steenrod operation to an invariant produces a potentially new invariant. The complete Steenrod operator $\mathcal{P}(t):S_{m}\to S_{m}[t]$ is the algebra homomorphism determined by $\mathcal{P}(t)(v)=v+v^{q}t$ for $v$ homogeneous of degree one. Since the map is linear in degree one, $\mathcal{P}(t)$ is well-defined. For $f$ homogeneous of degree $d$, the Steenrod operations $\mathcal{P}^{i}(f)$ are defined by
$\mathcal{P}(t)(f)=\sum_{i=0}^{d}\mathcal{P}^{i}(f)t^{i}.$
Note that for $i>d$ or $i<0$, $\mathcal{P}^{i}(f)=0$. It is clear that $\mathcal{P}^{0}(f)=f$ and $\mathcal{P}^{d}(f)=f^{q}$, i.e., the stability property is satisfied. The Steenrod operations satisfy the Cartan identity: for $f_{1},f_{2}\in S_{m}$
$\mathcal{P}^{i}(f_{1}f_{2})=\sum_{j=0}^{i}\mathcal{P}^{j}(f_{1})\mathcal{P}^{i-j}(f_{2}).$
The Steenrod operations also satisfy the Adem relations: for $i<qj$
$\mathcal{P}^{i}\mathcal{P}^{j}=\sum_{k}(-1)^{i+k}\binom{(q-1)(j-k)-1}{i-qk}\mathcal{P}^{i+j-k}\mathcal{P}^{k}.$
We can extend the action of ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$ to $S_{m}[t]$ by taking $tg=t$ for all $g$. Using this action, since taking a $q^{th}$ power is linear in $S_{m}$, we see that $\mathcal{P}(t)$ commutes with the ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$-action and $\mathcal{P}^{i}g=g\mathcal{P}^{i}$ for all $i$.
The following lemma is a consequence of the Cartan identity.
Lemma 3.3.
For $f\in S_{m}$, we have $\mathcal{P}^{i}(f^{q})=0$ unless $q$ divides $i$, in which case $\mathcal{P}^{i}(f^{q})=(\mathcal{P}^{i/q}(f))^{q}$.
Lemma 3.4.
Suppose $v,f\in S_{m}$ with $v$ homogeneous of degree one. Then $v$ divides $\mathcal{P}^{i}(vf)$ for all $i$.
Proof.
By definition, $v$ divides $\mathcal{P}^{j}(v)$. Therefore, using the Cartan identity, $v$ divides $\mathcal{P}^{i}(vf)$. ∎
It is an immediate consequence of the lemma that if $f$ is a product of linear forms, $f$ divides $\mathcal{P}^{i}(f)$.
Corollary 3.5.
(a) $\mathcal{P}(t)(\xi_{0})=\xi_{0}+\xi_{1}t+\xi_{0}^{q}t^{2}$.
(b) $\mathcal{P}(t)(\xi_{1})=\xi_{1}+2\xi_{0}^{q}t+\xi_{2}t^{q}+\xi_{1}^{q}t^{q+1}$.
(c) For $i>1$, $\mathcal{P}(t)(\xi_{i})=\xi_{i}+\xi_{i-1}^{q}t+\xi_{i+1}t^{q^{i}}+\xi_{i}^{q}t^{q^{i}+1}$.
Lemma 3.6.
Suppose $b\in S_{m}$ is homogeneous of degree $j>0$ and $a\in S_{m}$ is homogeneous of degree $i>0$.
(a) $\mathcal{P}^{i+j-1}(ab)=a^{q}\mathcal{P}^{j-1}(b)+b^{q}\mathcal{P}^{i-1}(a)$.
(b) $\mathcal{P}^{j+1}(\xi_{0}b)=\xi_{1}b^{q}+\xi_{0}^{q}\mathcal{P}^{j-1}(b)$.
(c) $\mathcal{P}^{j+3}(\xi_{0}^{2}b)=2\xi_{0}^{q}\xi_{1}b^{q}+\xi_{0}^{2q}\mathcal{P}^{j-1}(b)$.
(d) $\mathcal{P}^{2q+j-1}(\xi_{0}^{q}b)=\xi_{0}^{q^{2}}\mathcal{P}^{j-1}b$.
Proof.
Part (a) follows from the Cartan identity and the stability condition. Parts (b), (c) and (d) follow from (a) and Corollary 3.5. ∎
3.7. New Invariants
Let $P_{m}$ denote the intersection of $G_{m}$ with the upper triangular unipotent subgroup of ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$. $P_{m}$ is a choice of Sylow $p$-subgroup for $G_{m}$. For $v\in S_{m}$ homogenous of degree one, define $N(v)$ to be the orbit product $v$ over $P_{m}$. Since $x_{1}\in S_{m}^{P_{m}}$, $N(x_{1})=x_{1}$. We will see in Section 5 that $\deg(N(x_{i}))=q^{i-1}$ and $\deg(N(y_{i}))=q^{n-i-1}$. Define
$u_{m}:=\prod_{i=1}^{m}N(y_{i})N(x_{i})$
and observe that $\deg(u_{m})=(q^{m-1}+1)(1+q+\cdots+q^{m-1})$.
Remark 3.8.
We observe that each linear form that occurs as a factor of $u_{m}$ is isotropic under the induced symmetric bilinear form, $\beta^{*}$, on $V^{*}$. Furthermore every isotropic linear form is a scalar multiple of one of the factors of $u_{m}$.
Since $u_{m}$ is a product of linear forms, Lemma 3.4 implies that $u_{m}$ divides $\mathcal{P}^{j}(u_{m})$ in $S_{m}$. Define $e(i,m):=\sum_{j=1}^{i}q^{n-1-j}$ with $n=2m$ and, for $1\leq i\leq m$, define $d_{i,m}\in S_{m}$ by
$u_{m}d_{i,m}=\mathcal{P}^{e(i,m)}(u_{m}).$
For convenience, define $d_{0,m}=1$. Observe that $e(i,m)=q^{n-2}+qe(i-1,m-1)$.
Let $\psi:S_{m}\to S_{m}[t]$ denote the algebra homomorphism determined by $\psi(v)=v^{q}-vt^{q-1}$ for $v$ homogeneous of degree one. Since the map is linear in degree one, $\psi$ is well-defined. By comparing $\psi$ and the complete Steenrod operator $\mathcal{P}(t)$, we see that, for $f$ homogeneous of degree $d$,
$\psi(f)=\sum_{\ell=0}^{d}\mathcal{P}^{d-\ell}(f)(-t^{q-1})^{\ell}.$
Since $u_{m}$ divides $\mathcal{P}^{i}(u_{m})$ in $S_{m}$, we see that $\psi(u_{m})/u_{m}\in S_{m}[t]$ is monic with degree $(q-1)\deg(u_{m})$ as a polynomial in $t$.
Theorem 3.9.
$\psi(u_{m})/u_{m}$ is the minimal polynomial of $x_{1}$ over the field $\mathcal{F}(S_{m}^{G_{m}})$. In particular, $d_{i,m}\in S_{m}^{G_{m}}$ for $i=1,\ldots,m$.
Proof.
If we identify $G_{m-1}$ with the pointwise stabiliser of $\{y_{1},x_{1}\}$ in $G_{m}$, then the pointwise stabiliser of $x_{1}$ coincides with $G_{m-1}H$, where $H$ is the associated Hook group (see Section 5.1). From this we see that the size the orbit of $x_{1}$ is
$\frac{|{{\text{O}^{+}_{{n}}(\mathbb{F}_{q})}}|}{|{{\text{O}^{+}_{{n-2}}(\mathbb{F}_{q})}}||H|}=\frac{(q^{m}-1)(q^{2m-2})}{q^{m-1}-1}=(q^{m}-1)(q^{m-1}+1)=(q-1)\deg(u_{m})$
and this orbit consists of the non-zero scalar multiples of the linear factors of $u_{m}$. If $v$ is a linear factor of $u_{m}$, then the roots of $\psi(v)$ are the non-zero scalar multiples of $v$. From this, using the fact that $\psi(u_{m})/u_{m}$ is monic of degree $(q-1)\deg(u_{m})$, we conclude that $\psi(u_{m})/u_{m}=\prod\{t-x_{1}g\mid g\in G_{m}\}\in S_{m}^{G_{m}}$. Therefore, $\psi(u_{m})/u_{m}$ is the minimal polynomial of $x_{1}$ over $\mathcal{F}(S_{m}^{G_{m}})$ and, since the $d_{i,m}$ are coefficients of $\psi(u_{m})/u_{m}$ up to sign, we have $d_{i,m}\in S_{m}^{G_{m}}$. ∎
Lemma 3.10.
Using the lexicographic order on $S_{m}$,
$\operatorname{LT}(u_{m})=\prod_{j=1}^{m}y_{j}^{q^{n-j-1}}x_{i}^{q^{j-1}}$
and $\operatorname{LT}(d_{1,m})=y_{1}^{q^{n-1}-q^{n-2}}$.
Proof.
The expression for $\operatorname{LT}(u_{m})$ follows directly from the definition of $u_{m}$. We apply $\mathcal{P}^{q^{n-2}}$ to $u_{m}$ to get $u_{m}d_{1,m}$. Using the Cartan identity this gives
$u_{m}d_{1,m}=\mathcal{P}^{q^{n-2}}\left(N(y_{1})\right)x_{1}\prod_{i=2}^{m}N(y_{i})N(x_{i})+F$
where $F$ is smaller than $\operatorname{LT}(\mathcal{P}^{q^{n-2}}(N(y_{1}))=y_{1}^{q^{n-1}}$ in the lexicographic order. Dividing by $u_{m}$ gives $\operatorname{LT}(d_{1,m})=y_{1}^{q^{n-1}-q^{n-2}}$. ∎
Define :=$\mathcal{H}:=\{\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{m,m}\}$.
Theorem 3.11.
$\mathcal{H}$ is a homogeneous system of parameters.
Proof.
We will show that the variety in $\overline{V}=\overline{\mathbb{F}}_{q}\otimes V$ cut out by the ideal generated by $\mathcal{H}$ is $\{\underline{0}\}$. Using Theorem 3.1,
$\mathcal{V}(\xi_{0},\ldots,\xi_{m-1})=\bigcup_{g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}g\mathcal{V}(x_{1},x_{2},\ldots,x_{m}).$
For $v\in\mathcal{V}(\xi_{0},\ldots,\xi_{m-1})$, choose $g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ so that $gv\in\mathcal{V}(x_{1},x_{2},\ldots,x_{m})$. From Theorem 3.9, $d_{i,m}\in S_{m}^{G_{m}}$. Therefore $d_{i,m}(gv)=d_{i,m}(v)$. Hence to show $\mathcal{V}(\mathcal{H})=\{\underline{0}\}$, it is sufficient to show $\mathcal{V}(x_{1},\ldots,x_{m},d_{1,m},\ldots,d_{m,m})=\{\underline{0}\}$. To do this we consider $d_{i,m}$ modulo the ideal $I:=\langle x_{1},x_{2},\ldots,x_{m}\rangle$.
Define $\bar{d}_{i}\in\mathbb{F}_{q}[y_{1},\ldots,y_{m}]$ by $\bar{d}_{i}\equiv_{I}d_{i,m}$. We will show that $\{\bar{d}_{1},\ldots,\bar{d}_{m}\}$ is an hsop in $\mathbb{F}_{q}[y_{1},\ldots,y_{m}]$. Writing $d_{1,m}$ in terms of the generators of the Borel invariants (see Section 5.18), shows that $d_{1,m}$ is congruent to an element of the ring $\mathbb{F}_{q}[N(y_{1})^{q-1},\ldots,N(y_{m})^{q-1}]$. Define $W_{i}:={\rm Span}_{\mathbb{F}_{q}}\{y_{i},\ldots,y_{m}\}$ and observe that
$N(y_{i})\equiv_{I}\prod_{w\in W_{i+1}}(y_{i}+w)^{q^{m-1}}=\prod_{w\in W_{i+1}}(y_{i}^{q^{m-1}}+w^{q^{m-1}})$
(for $i=m$, $W_{m+1}=0$ and $N(y_{m})\equiv_{I}y_{m}^{q^{m-1}}$). Therefore $\bar{d}_{1}\in\mathbb{F}_{q}[y_{1}^{q^{m-1}},\ldots,y_{m}^{q^{m-1}}]$. Observe that taking $g$ to $g\oplus(g^{-1})^{T}$ gives an embedding of ${{\text{GL}_{{m}}(\mathbb{F}_{q})}}$ into ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. Using this embedding and the fact that $d_{1,m}\in S_{m}^{G_{m}}$, we have
$\bar{d}_{1}\in\mathbb{F}_{q}[y_{1}^{q^{m-1}},\ldots,y_{m}^{q^{m-1}}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}}\ .$
By comparing lead terms (see Lemma 3.10), we see that
$\bar{d}_{1}=d_{1}(y_{1}^{q^{m-1}},\ldots,y_{m}^{q^{m-1}})=d_{1}(W_{1})^{q^{m-1}},$
where $d_{1}(W_{1})$ is the first Dickson invariant for $W_{1}$. For a definition and background material on the Dickson invariants see [24] or [12] or [5, §3.3]. We claim that
$d_{i,m}\equiv_{I}\pm d_{i}(W_{1})^{q^{m-1}}$
for all $i$. From this, it follows that $\{\bar{d}_{1},\ldots,\bar{d}_{m}\}$ is an hsop for $\mathbb{F}_{q}[y_{1},\ldots,y_{m}]$, as required.
To see that $d_{i,m}\equiv_{I}\pm d_{i}(W_{1})^{q^{m-1}}$, we look at $\psi(u_{m})/u_{m}$ modulo $I$:
$\psi(u_{m})/u_{m}=\prod\{t-x_{1}g\mid g\in G_{m}\}\equiv_{I}t^{q^{m}-1}\left(\prod_{w\in W_{1}\setminus\{0\}}(t-w)\right)^{q^{m-1}}.$
On the right-hand side of this equivalence, $d_{i}(W_{1})^{q^{m-1}}$ is the coefficient of
$t^{q^{m}-1}(t^{q^{m-i}-1})^{q^{m-1}}=t^{q^{2m-i-1}+q^{m}-q^{m-1}-1}.$
On the left-hand side of the equivalence, $d_{i,m}$ is the coefficient of $(-t^{q-1})^{\ell}$ with
$\ell(q-1)$ $=$ $(q^{m-1}+1)(q^{m}-1)-(q-1)e(i,m)$
$=$ $q^{2m-1}+q^{m}-q^{m-1}-1-(q^{n-1}-q^{n-i-1})$
$=$ $q^{n-i-1}+q^{m}-q^{m-1}-1$
from which we conclude $d_{i,m}\equiv_{I}\pm d_{i}(W_{1})^{q^{m-1}}$, where the parity is determined by $(-1)^{\ell}=(-1)^{i+1}$. ∎
Remark 3.12.
We observe that the orthogonal invariants $d_{i,m}$ are in certain ways analogous to the Dickson invariants $d_{i}(V^{*})$ which generate $\mathbb{F}_{q}[V]^{\text{GL}(V)}$. In particular, in both cases these invariants are certain elementary symmetric functions in the orbit of a linear form. Furthermore (see Theorem 8.1) these two families of invariants may each be viewed as arising from a single invariant under the action of the Steenrod algebra. In the analogous calculation for the invariants of the symplectic group the $d_{i}(V^{*})$ play the role of the $d_{i,m}$.
Remark 3.13.
It follows from the proof of Theorem 3.11 that using the grevlex order on $S_{m}$,
$\operatorname{LT}(d_{i,m})=\prod_{j=1}^{i}y_{j}^{q^{n-j}-q^{n-j-1}}.$
4. Formulating the Main Theorem
Define $R_{i}:=\mathbb{F}_{q}[T_{i},T_{i-1},\ldots,T_{0}]$ and let $\Phi_{i,m}:R_{i}\to S_{m}$ denote the algebra map defined by $\Phi_{i,m}(T_{j})=\xi_{j}$. If we grade $R_{i}$ by assigning $T_{j}$ degree $q^{j}+1$, then $\Phi_{i,m}$ preserves degree. Since $\{\xi_{0},\ldots,\xi_{2m-1}\}$ is algebraically independent in $S_{m}$, $\Phi_{i,m}$ is injective for $i<2m$.
Lemma 4.1.
The kernel of $\Phi_{2m,m}$ is a principal ideal. Furthermore, the generator is of the form $aT_{2m}-b$ with $a,b\in R_{2m-1}\setminus\{0\}$.
Proof.
The field of fractions of $S_{m}^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ is $\mathbb{F}_{q}(\xi_{0},\ldots,\xi_{2m-1})$. Write $\xi_{2m}=\overline{b}/\overline{a}$ with $\overline{a},\overline{b}\in\mathbb{F}_{q}[\xi_{0},\ldots,\xi_{2m-1}]=\operatorname{image}(\Phi_{2m-1,m})$. We may assume that $\overline{a}$ and $\overline{b}$ are relatively prime. Since $\Phi_{2m-1,m}$ is injective, there are unique $a,b\in R_{2m-1}$ with $\Phi_{2m-1,m}(a)=\overline{a}$ and $\Phi_{2m-1,m}(b)=\overline{b}$. Evaluating $\Phi_{2m,m}(aT_{2m}-b)$ gives $\overline{a}\xi_{2m}-\overline{b}=0$. Therefore $aT_{2m}-b$ lies in the kernel of $\Phi_{2m,m}$. For any $F\in R_{2m}$, for $\ell$ sufficiently large we can divide by $aT_{2m}-b$ to get $a^{\ell}F=(aT_{2m}-b)f+r$ with $f,r\in R_{2m-1}$. Applying $\Phi_{2m,m}$ gives
$\Phi_{2m,m}(a^{\ell}F)=\overline{a}^{\ell}\Phi_{2m,m}(F)=\Phi_{2m-1,m}(r).$
Since $\Phi_{2m-1,m}$ is injective we see that $a^{\ell}F$ is in $\ker(\Phi_{2m,m})$ if and only if $a^{\ell}F=(aT_{2m}-b)f$. Since $\overline{a}$ and $\overline{b}$ are relatively prime, $aT_{2m}-b$ is prime in $R_{2m}$. Hence $F\in\ker(\Phi_{2m,m})$ if and only if $F$ is divisible by $aT_{2m}-b$. ∎
Let $\overline{\Phi}_{i,m}$ denote the composition of $\Phi_{i,m}$ followed by projection onto $S_{m}/x_{1}S_{m}$. Let $\sigma$ denote the algebra monomorphism from $S_{m-1}$ to $S_{m}$ which takes $y_{j}$ to $y_{j+1}$ and $x_{j}$ to $x_{j+1}$ for $j=1,\ldots,m-1$. Let $\psi_{1}$ denote the algebra homomorphism from $S_{m}$ to $S_{m}$ which takes $w$ to $w^{q}-wx_{1}^{q-1}$ for $w$ homogeneous of degree one.
Lemma 4.2.
(a) $\ker(\psi_{1})=x_{1}S_{m}$ and $\psi_{1}\circ\psi_{1}(f)=\psi_{1}(f)^{q}$.
(b) $\ker(\overline{\Phi}_{i,m})=\ker(\Phi_{i,m-1})=\ker(\psi_{1}\circ\Phi_{i,m}).$
(c) If $f\in\ker(\overline{\Phi}_{i,m})$ then $u_{m}$ divides $\Phi_{i,m}(f)$.
Proof.
(a) Clearly $x_{1}S_{m}\subseteq\ker(\psi_{1})$. To see that $x_{1}S_{m}=\ker(\psi_{1})$, identify $S_{m}/x_{1}S_{m}$ with $\mathbb{F}_{q}[y_{1},y_{2},\ldots,x_{2}]$ and observe that $\psi_{1}$ maps this ring injectively to $\mathbb{F}_{q}[y_{1}^{q}-y_{1}x_{1}^{q-1},y_{2}^{q}-y_{2}x_{1}^{q-1},\ldots,x_{2}^{q}-x_{2}x_{1}^{q-1}]$. For $w\in S_{m}$ with $\deg(w)=1$, we have $\psi_{1}\circ\psi_{1}(w)=\psi_{1}(w^{q}-wx_{1}^{q-1})=\psi_{1}(w)^{q}$. The general result follows from the fact that taking the $q^{th}$ power is an algebra monomorphism on $S_{m}$.
(b) If we identify $S_{m}/x_{1}S_{m}$ with $\sigma(S_{m-1})[y_{1}]$, then $\overline{\Phi}_{i,m}=\sigma\circ\Phi_{i,m-1}$. Since $\sigma$ is a monomorphism, we see that $\ker(\overline{\Phi}_{i,m})=\ker(\Phi_{i,m-1})$. From (a) we have $\ker(\psi_{1})=x_{1}S_{m}$. Therefore $\ker(\overline{\Phi}_{i,m})=\ker(\psi_{1}\circ\Phi_{i,m})$ follows from the definition of $\overline{\Phi}_{i,m}$.
(c) If $f\in\ker(\overline{\Phi}_{i,m})$ then $x_{1}$ divides $\Phi_{i,m}(f)$. Since $\Phi_{i,m}(f)$ is invariant, if $x_{1}$ divides $\Phi_{i,m}(f)$, each element in the orbit of $x_{1}$ must divide $\Phi_{i,m}(f)$. The result follows from the fact that $u_{m}$ is a product of independent linear factors from the orbit of $x_{1}$
In the following, for $i<2m$, we identify $R_{i}$ and $\Phi_{i,m}(R_{i})$. Recall that $\psi:S_{m}\to S_{m}[t]$ is the algebra homomorphism determined by $\psi(v)=v^{q}-vt^{q-1}$ for $v$ homogeneous of degree one. Note that $\psi$ followed by evaluating $t$ at $x_{1}$ gives $\psi_{1}:S_{m}\to S_{m}$. Also note that $\psi(\xi_{0})=\xi_{0}^{q}-\xi_{1}t^{q-1}+\xi_{0}t^{2(q-1)}$, $\psi(\xi_{1})=\xi_{1}^{q}-\xi_{2}t^{q-1}-2\xi_{0}^{q}t^{q(q-1)}+\xi_{1}t^{(q+1)(q-1)}$, and $\psi(\xi_{i})=\xi_{i}^{q}-\xi_{i+1}t^{q-1}-\xi_{i-1}^{q}t^{q^{i}(q-1)}+\xi_{i}t^{(q^{i}+1)(q-1)}$ for $i>1$. Therefore $\psi$ restricts to a map from $R_{n-2}$ to $R_{n-1}[t]$.
For $f\in R_{n-1}$, define ${R\mbox{-}\deg}(f)$ to be the degree of $f$ as a polynomial in the variables $\xi_{0},\ldots,\xi_{n-1}$. We will use the corresponding weighted grevlex order on $R_{n-1}$.
Define $\nu:R_{n-1}\to\mathbb{Z}\cup\{\infty\}$ by $\nu(0)=\infty$ and, for $f\not=0$, $\nu(f)$ is the minimum value of ${R\mbox{-}\deg}$ on the non-zero terms of $f$. For example, $\nu(\xi_{1}^{q})=q$ and $\nu(\xi_{1}^{q}-\xi_{2}\xi_{1}^{(q-1)/2})=(q+1)/2$. Note that if $\nu(f)=\ell$, then $f\in(R_{n-1})_{+}^{\ell}\setminus(R_{n-1})_{+}^{\ell+1}$. It is easy to verify that for $f,h\in R_{n-1}$, we have $\nu(fh)=\nu(f)\nu(h)$ and $\nu(f+h)\geq{\rm min}\{\nu(f),\nu(h)\}$. It follows from [25, Ch IV, §9, Theorem 14], that defining $\nu(f/h)=\nu(f)-\nu(h)$ extends $\nu$ to a discrete valuation on $\mathcal{F}(R_{n-1})$. Note that the associated Discrete Valuation Ring, $\{a\in\mathcal{F}(R_{n-1})\mid\nu(a)\geq 0\}$, properly contains $R_{n-1}$, i.e., $R_{n-1}$ is not the associated Discrete Valuation Ring.
The following is a consequence of Corollary 3.5.
Lemma 4.3.
For $f\in R_{n-1}$, we have $\nu(\mathcal{P}^{i}(f))\geq\nu(f)$.
For $f\in R_{n-1}$, define ${S\mbox{-}\deg}(f)$ to be the degree of $f$ as an element of $S_{m}$. For example, ${S\mbox{-}\deg}(\xi_{i})=q^{i}+1$ while ${R\mbox{-}\deg}(\xi_{i})=1$.
Lemma 4.4.
Suppose $\beta$ is a monomial in $R_{n-1}$ with ${S\mbox{-}\deg}(\beta)={S\mbox{-}\deg}(\xi_{i}^{b}\xi_{i+1}^{a})$. If $\nu(\beta)>\nu(\xi_{i}^{b}\xi_{i+1}^{a})$ then $\operatorname{LT}(\beta)<\operatorname{LT}(\xi_{i}^{b}\xi_{i+1}^{a})$ using the weighted grevlex order.
Proof.
It is sufficient to show that $\beta$ is in the ideal $\langle\xi_{0},\ldots,\xi_{i-1},\xi_{i}^{b+1}\rangle\subset R_{n-1}$. Suppose, by way of contradiction, that this is not the case. Since $\nu(\beta)>b+a$, $\beta$ has at least $a+b+1$ factors. Since $\beta$ is not in the ideal $\langle\xi_{0},\ldots,\xi_{i-1},\xi_{i}^{b+1}\rangle$, at least $a+1$ of these factors must be greater than $\xi_{i}$. The smallest weighted degree of such a monomial is
$b(q^{i}+1)+(a+1)(q^{i+1}+1)>b(q^{i}+1)+a(q^{i+1}+1)={S\mbox{-}\deg}(\xi_{i}^{b}\xi_{i+1}^{a}),$
contradicting the hypothesis that ${S\mbox{-}\deg}(\beta)={S\mbox{-}\deg}(\xi_{i}^{b}\xi_{i+1}^{a})$. ∎
Define an $m\times(m+1)$ matrix
$M_{m}:=\begin{pmatrix}\xi_{n-1}&\xi_{n-2}&&\cdots&\xi_{m-1}\\ \xi_{n-2}^{q}&\xi_{n-3}^{q}&&\cdots&\xi_{m-2}^{q}\\ &&\vdots&&\\ \xi_{m}^{q^{m-1}}&\xi_{m-1}^{q^{m-1}}&&\cdots&\xi_{0}^{q^{m-1}}\end{pmatrix}$
and let $M(i,m)$ denote the minor formed by removing column $i+1$ from $M_{m}$. Note that $\nu(M(i,m))=1+q+\cdots+q^{m-1}$. It follows from Lemma 6.1 that $\nu(u_{2})=q+1$ and $u_{2}=M(0,2)+\delta_{0,2}$ with $\nu(\delta_{0,2})=2q$. It follows from Lemma 6.4 that $u_{2}d_{1,2}=M(1,2)+\delta_{1,2}$ with $\nu(\delta_{1,2})=2q$ and $u_{2}d_{2,2}=M(2,2)+\delta_{2,2}$ with $\nu(\delta_{2,2})=q^{2}+1$.
Lemma 4.5.
$\nu(\mathcal{P}^{e(i,m)}(M(0,m))-M(i,m))>1+q+\cdots+q^{m-1}$.
Proof.
Let $m_{j,k}$ denote the row $j$ column $k$ entry of $M_{m}$. Then
$M(0,m)=\sum_{\sigma\in\Sigma_{m}}\operatorname{sgn}(\sigma)\left(\prod_{k=1}^{m}m_{\sigma(k),k+1}\right)$
and, for $0<i\leq m$,
$M(i,m)=\sum_{\sigma\in\Sigma_{m}}\operatorname{sgn}(\sigma)\left(\prod_{k=1}^{i}m_{\sigma(k),k}\right)\left(\prod_{k=i+1}^{m}m_{\sigma(k),k+1}\right).$
Note that $\nu(m_{j,k})=q^{j-1}$. For $k>1$, it follows from Lemma 3.5 that $\mathcal{P}^{q^{n-k}}(m_{j,k})=m_{j,k-1}$, and $\nu(\mathcal{P}^{\ell}(m_{j,k}))>q^{j-1}$ unless $\ell=0$ or $\ell=q^{n-k}$. Therefore, since $e(i,m)=q^{n-2}+q^{n-3}+\cdots+q^{n-i-1}$, we have
$\mathcal{P}^{e(i,m)}\left(\prod_{k=1}^{m}m_{\sigma(k),k+1}\right)=\left(\prod_{k=1}^{i}m_{\sigma(k),k}\right)\left(\prod_{k=i+1}^{m}m_{\sigma(k),k+1}\right)+\delta$
with $\nu(\delta)>1+q+\cdots+q^{m-1}$. Hence $\nu(\mathcal{P}^{e(i,m)}(M(0,m))-M(i,m))>1+q+\cdots+q^{m-1}$. ∎
We will show below that $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$ with $\nu(\delta_{i,m})>\nu(M(i,m))=1+q+\cdots+q^{m-1}$.
Finally we are ready to formulate the main theorem. Recall from Theorem 3.11 the homogeneous system of parameters
=$\mathcal{H}=\{\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{m,m}\}$
and let $\mathcal{B}$ denote the set of monomial factors of $\prod_{i=m}^{n-2}\xi_{i}^{q^{n-i}-1}$.
Theorem 4.6.
Suppose $m>1$.
(a) $u_{m}\in R_{n-2}$, $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$, $u_{m}d_{i,m}\in R_{n-1}$ and $d_{i,m}\in S_{m}^{G_{m}}$. Furthermore, interpreting $u_{m-1}$ and $u_{m-1}d_{i,m-1}$ as elements of $R_{n-3}\subset S_{m}^{G_{m}}$, we have
$u_{m}=(\xi_{n-2}+c_{n-2,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(u_{m-1}d_{i,m-1})^{q}$
with $c_{j,m}\in R_{j-1}$, $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$, $\mathcal{P}^{i}(u_{m})=0$ for $0<i<q^{m-1}$ and $\mathcal{P}^{i}(u_{m})/u_{m}\in R_{n-2}$ for $q^{m-1}\leq i<q^{n-2}$. For $i=\ell q^{n-2}+r$ with $0\leq\ell<q$ and $0\leq r<q^{n-2}$, $u_{m}^{\ell+1}\mathcal{P}^{i}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$. If $\ell=q-1$ and $i\not\equiv_{(q)}0$ then $u_{m}^{q}\mathcal{P}^{i}(d_{j,m})$ has $\xi_{n-1}$-degree at most $q-1$.
(b) Using the weighted grevlex order on $R_{n-1}$,
$\operatorname{LT}(u_{m})=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1}+q^{m-2}+\cdots+1}$
and $\operatorname{LT}(u_{m}d_{i,m})=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1-i}+\cdots+1}\xi_{m}^{q^{m-1}+\cdots+q^{m-i}}$. Furthermore
$\nu(u_{m})=\nu(u_{m}d_{i,m})=1+q+\cdots+q^{m-1},$
$u_{m}=M(0,m)+\delta_{0,m}$ with $\delta_{0,m}\in R_{n-2}$, $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$ with $\delta_{i,m}\in R_{n-1}$ and, for $0\leq i\leq m$, $\nu(\delta_{i,m})>1+q+\cdots+q^{m-1}$.
(c) $\mathcal{H}$ is a homogeneous system of parameters.
(d) There exists $c_{n-1,m}\in R_{n-2}$ such that
$\xi_{n-1}=\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-1-i}+c_{n-1-i,m})d_{i,m}\right)-c_{n-1,m}.$
(e) For $11\leq i\leq m-$ 1 there exist $\gamma_{k,m}^{(i)}\in R_{n-2}$ with $\nu(\gamma_{k,m}^{(i)})>q^{i}$ such that
$\xi_{n-1-i}^{q^{i}}=\left(\sum_{j=1}^{m}(-1)^{j+1}(\xi_{n-1-i-j}^{q^{i}}+\gamma_{n-1-i-j,m}^{(i)})d_{j,m}\right)-\gamma_{n-1-i,m}^{(i)}.$
(f) $S_{m}^{G_{m}}$ is generated by $\mathcal{H}\cup\{\xi_{m},\ldots,\xi_{n-2}\}$. Furthermore, $S_{m}^{G_{m}}$ is the free $\mathbb{F}_{q}[\mathcal{H}]$-module with basis $\mathcal{B}$ and is a complete intersection with the relations given in (e).
Remark 4.7.
The rank of $S_{m}^{G_{m}}$ as a free $\mathbb{F}_{q}[\mathcal{H}]$-module is $q^{m(m-1)/2}$.
Remark 4.8.
Note that the relations in the associated graded algebra are given by
$\xi_{n-1-i}^{q^{i}}=\sum_{j=1}^{m}(-1)^{j+1}\xi_{n-1-i-j}^{q^{i}}d_{j,m}\quad\text{ for }i=1,2,\dots,m-1.$
This formula is also true for $i=0$.
Corollary 4.9.
$S_{m}^{G_{m}}$ is minimally generated by
$\{\xi_{0},\xi_{1},\dots,\xi_{n-2},d_{1,m},d_{2,m},\dots,d_{m,m}\}\ .$
Proof.
The invariants $\xi_{0},\xi_{1},\dots,\xi_{n-2}$ are algebraically independent and of degree less than $\deg(d_{i,m})$ for all $i$. Hence none of these $\xi_{i}$ is redundant. But Theorem 3.1 shows that at least $m$ other invariants are required to generate $S_{m}^{G_{m}}$. ∎
5. Sylow Invariants
The order of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ is $2q^{m(m-1)}(q^{m}-1)\prod_{j=1}^{m-1}(q^{2j}-1)$, see [23, page 141]. Therefore a Sylow $p$-subgroup has order $q^{m(m-1)}$.
The field of fractions for the Sylow invariants was calculated in [14] and the ring of invariants for the $m=2$ case was calculated in [15].
Let $H$ denote the Hook subgroup of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ consisting of unipotent matrices which differ from the identity only in the first row and the last column. A direct calculation shows that for $h\in H$ we have
$y_{1}\cdot h=y_{1}+\sum_{i=2}^{m}(b_{i}y_{i}+a_{i}x_{i}-a_{i}b_{i}x_{1}),$
$x_{1}\cdot h=x_{1}$, and for $j>1$, $y_{j}\cdot h=y_{j}-a_{j}x_{1}$, $x_{j}\cdot h=x_{j}-b_{j}x_{1}$ where $a_{i},b_{i}\in\mathbb{F}_{q}$. Hence the order of $H$ is $q^{2m-2}$. Note that $H$ is a normal subgroup of $P_{m}$. Identify $P_{m-1}$ with pointwise stabiliser of $\{y_{1},x_{1}\}$ in $P_{m}$. Then $P_{m}=P_{m-1}H$. (Note that $(m-1)(m-2)+(2m-2)=m(m-1)$.) Therefore $S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}$. We will calculate $S_{m}^{P_{m}}$ by first computing the Hook invariants.
5.1. Computing the Hook Invariants
Consider the subalgebra
$Q:=\mathbb{F}_{q}[y_{2},\ldots,y_{m},x_{m},\ldots,x_{1}]\subset S_{m}.$
For $i>1$, define $Y_{i}:=y_{i}^{q}-y_{i}x_{1}^{q-1}$ and $X_{i}=x_{i}^{q}-x_{i}x_{1}^{q-1}$. It is easy to see that
$Q^{H}:=\mathbb{F}_{q}[x_{1}][Y_{i},X_{i}\mid i=2,\ldots,m].$
Since $\xi_{0}$ has degree $1$ in $y_{1}$, we have $S_{m}^{H}[x_{1}^{-1}]=Q^{H}[\xi_{0},x_{1}^{-1}]$ (see [3, Theorem 2.4]).
A straight-forward calculation gives
(1) $\xi_{0}^{q}-\xi_{1}x_{1}^{q-1}+\xi_{0}x_{1}^{2q-2}=\sum_{i=2}^{m}X_{i}Y_{i}=\psi_{1}(\xi_{0}).$
For $j>0$, further calculation, or the application of Steenrod operations, gives
(2) $\xi_{j}^{q}-\xi_{j+1}x_{1}^{q-1}=\xi_{1}x_{1}^{q^{j+1}-1}-2\xi_{0}x_{1}^{q^{j+1}+q-2}+\psi_{1}(\xi_{j}).$
Observe that the orbit product of $y_{1}$ over $H$ coincides with the orbit product of $y_{1}$ over $P_{m}$.
Lemma 5.2.
$x_{1}N(y_{1})-\xi_{n-2}\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$
We postpone the proof of the lemma to the end of the section.
Define
$A:=Q^{H}[N(y_{1}),\xi_{j}\mid j=0,\ldots,n-3]$
and $\mathcal{W}:=\{X_{i},Y_{i}\mid i=2,\ldots,m\}$.
Lemma 5.3.
The set
$\mathcal{W}\cup\{N(y_{1}),x_{1},\xi_{j}\mid j=0,\ldots,n-3\}$
is a Khovanskii basis for $A$ using the lexicographic order. Furthermore, $A$ is a complete intersection with relations given by the subductions of the tête-a-têtes $\xi_{n-3}^{q}-x_{1}^{q}N(y_{1})$ and $\xi_{j}^{q}-x_{1}^{q-1}\xi_{j+1}$ for $j=0,\ldots,n-4$.
Proof.
The tête-a-têtes $\xi_{j}^{q}-\xi_{j+1}x_{1}^{q-1}$ for $j=1,\ldots,n-4$ subduct to zero using Equations 1 and 2. The subduction of the tête-a-tête $x_{1}^{q}N(y_{1})-\xi_{n-3}^{q}$ is constructed using Equation 2 with $j=n-3$ and Lemma 5.2 to substitute for $x_{1}^{q-1}\xi_{n-2}$. ∎
Theorem 5.4.
$S_{m}^{H}=A$.
Proof.
Since $Q^{H}[N(y_{1})]$ is generated by a homogeneous system of parameters and $S_{m}^{H}[x_{1}^{-1}]=Q^{H}[\xi_{0},x_{1}^{-1}]$, it is sufficient to prove that the ring $A$ is integrally closed in its field of fractions. We do this by proving $x_{1}$ is prime in $A$ which is sufficient by [19, Theorem 20.2].
It follows from Lemma 5.3 that $A$ is Cohen-Macaulay and has rank $q^{n-2}$ over $Q^{H}[N(y_{1})]$ with a basis given by the monomial factors of $(\xi_{0}\cdots\xi_{n-3})^{q-1}$. We will show that $x_{1}$ is prime in $A$ by proving that $A/x_{1}A$ is an integral domain.
Recall that $\psi_{1}:S_{m}\to S_{m}$ is the algebra homomorphism determined by $\psi_{1}(w)=w^{q}-wx_{1}^{q-1}$ for $w$ homogeneous of degree one. Clearly $x_{1}S_{m}\subset\ker(\psi_{1})$. To see that $x_{1}S_{m}=\ker(\psi_{1})$, identify $S_{m}/x_{1}S_{m}$ with $\mathbb{F}_{q}[y_{1},\ldots,x_{2}]$ and observe that $\psi_{1}$ maps this ring injectively to $\mathbb{F}_{q}[y_{1}^{q}-y_{1}x_{1}^{q-1},Y_{2},\ldots,X_{2}]$.
Let $\overline{\psi_{1}}$ denote the algebra map from $A/x_{1}A$ to $S_{m}$ induced by $\psi_{1}$. If $\overline{\psi_{1}}$ is injective, then $A/x_{1}A$ is an integral domain, proving $x_{1}A$ is a prime ideal and $A$ is normal.
Define $\overline{Q}=\mathbb{F}_{q}[N(y_{1}),Y_{2},\ldots,X_{2}]$, i.e., the subalgebra generated by $\{N(y_{1})\}\cup\mathcal{W}$. Let $\mathcal{B}_{H}$ denote the set of monomial factors of $(\xi_{0}\cdots\xi_{n-3})^{q-1}$. Using the $Q^{H}$-module structure of $A$, we see that $A/x_{1}A$ is a free $\overline{Q}$-module with a basis $\mathcal{B}_{H}$. By identifying $\overline{Q}$ with the submodule of $A/x_{1}A$ generated by $1$ we see that $A/x_{1}A$ is a complete intersection with relations $\xi_{i}^{q}=\psi_{1}(\xi_{i})$ for $i=0,\ldots,n-3$. Observe that for $w\in S_{m}$ with $\deg(w)=1$, $\psi_{1}\circ\psi_{1}(w)=\psi_{1}(w)^{q}$. Therefore $\psi_{1}(A)$ is a module over $\mathbb{F}_{q}[\psi_{1}(N(y_{1})),Y_{2}^{q},\dots,X_{2}^{q}]$ with module generators $\{\psi_{1}(\beta)\mid\beta\in\mathcal{B}_{H}\}$. To prove that $\overline{\psi_{1}}$ is injective, it is sufficient to show that $\overline{\psi_{1}}(A/x_{1}A)=\psi_{1}(A)$ is a free module over $\mathbb{F}_{q}[\psi_{1}(N(y_{1})),Y_{2}^{q},\dots,X_{2}^{q}]$ with module generators $\{\psi_{1}(\beta)\mid\beta\in\mathcal{B}_{H}\}$. Note that for $\beta\in\mathcal{B}_{H}$, we have $\psi_{1}(\beta)\in\mathbb{F}_{q}[Y_{2}^{q},\dots,X_{2}^{q}]$. Therefore, it is sufficient to show that the $\mathbb{F}_{q}[Y_{2}^{q},\dots,X_{2}^{q}]$-module generated by $\{\psi_{1}(\beta)\mid\beta\in\mathcal{B}_{H}\}$ is free. After relabelling and a reindexing which takes $n-2$ to $n$, this is equivalent to showing that the $\mathbb{F}_{q}[y_{1}^{q},\dots,x_{1}^{q}]$-module generated by $\{\xi_{0}^{e_{0}}\cdots\xi_{n-1}^{e_{n-1}}\mid e_{i}<q\}\subset S_{m}$ is free of rank $q^{n}$.
Let $G$ denote a finite orthogonal group with defining representation $V$. We know that the invariant field $\mathbb{F}_{q}(V)^{G}$ is generated by $\{\xi_{0},\ldots,\xi_{n-1}\}$. Since the extension $\mathbb{F}_{q}(\xi_{0},\ldots,\xi_{n-1})=\mathbb{F}_{q}(V)^{G}\subset\mathbb{F}_{q}(V)$ is a Galois extension, for any intermediate extension $\mathbb{F}_{q}(V)^{G}\subset E\subset\mathbb{F}_{q}(V)$, we have $E=\mathbb{F}_{q}(V)^{L}$ for some subgroup $L\leq G$. Consider $E=\mathbb{F}_{q}(V)^{G}[y_{1}^{q},...,x_{1}^{q}]$. Since the pointwise stabiliser of $\{y_{1}^{q},...,x_{1}^{q}\}$ in ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$ is $\{1\}$, we have $\mathbb{F}_{q}(V)^{G}[y_{1}^{q},...,x_{1}^{q}]=\mathbb{F}_{q}(V)$. On the other hand, $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})\subset\mathbb{F}_{q}(V)$ is a field extension of degree $q^{n}$. Construct this extension by iteratively adjoining the $\xi_{i}$. Since $\xi_{i}$ is a root of $t^{q}-\xi_{i}^{q}\in\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})[t]$, at each step the extension has degree at most $q$. To complete the construction in $n$ steps, we need each step to be an extension of degree $q$. We can adjoin the $\xi_{i}$ in any order. Therefore $t^{q}-\xi_{i}^{q}$ is irreducible over $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})[\xi_{j}\mid j\not=i]$. Hence $\{\xi_{0}^{e_{0}}\cdots\xi_{n-1}^{e_{n-1}}\mid e_{i}<q\}$ is a basis for $\mathbb{F}_{q}(V)$ as a vector space over $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})$. This means that $\{\xi_{0}^{e_{0}}\cdots\xi_{n-1}^{e_{n-1}}\mid e_{i}<q\}$ is linearly independent over $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})$ and generates a free $\mathbb{F}_{q}[y_{1}^{q},...,x_{1}^{q}]$-module of rank $q^{n}$, completing the proof. ∎
The rest of Section 5.1 is devoted to the proof of Lemma 5.2. For a non-zero element $h\in S=S_{m}$ we denote by $\nu_{1}(h)$ the greatest integer $n$ such that $h\in(x_{1})^{n}S$, i.e., $x_{1}^{\nu_{1}(h)}$ divides $h$ in $S$ but $x_{1}^{\nu_{1}(h)+1}$ does not. Let $d$ be a non-negative integer. We write $||d||_{q}=||d||$ to denote the digit sum in base $q$, i.e., if $d=\sum_{i=0}^{r}d_{i}q^{i}$ with $0\leq d_{i}<q$ for all $i$ then $||d||_{q}=\sum_{i=0}^{r}d_{i}$. We will say that a polynomial $f=f(y_{1})=\sum_{k}c_{k}y_{1}^{k}\in S$ with each $c_{k}\in Q$ is compliant (with respect to $x_{1}$) if $\nu_{1}(c_{k})\geq||k||-1$ for all $k\geq 1$. We say that $f$ is strongly compliant if $\nu_{1}(c_{k})\geq||k||$ for all $k\geq 1$. Observe that $\xi_{i}$ is strongly compliant for all $i$ and that the set of strongly compliant polynomials is closed under addition. Furthermore, since $||k+\ell||\leq||k||+||\ell||$, the set of strongly compliant polynomials is closed under multiplication.
Lemma 5.5.
If $f\in S^{H}$ is strongly compliant and $\deg_{y_{1}}(f)<q^{n-2}$, then $f\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$.
Proof.
The proof is by induction on $d:=\deg_{y_{1}}(f)$. For $d=0$, $f=c_{0}\in Q^{H}$. Suppose the result is true for all strongly compliant invariants with $0<d<\ell<q^{n-2}$.
Consider $f=\sum_{k=0}^{\ell}c_{k}y_{1}^{k}\in S^{H}$ with $c_{k}\in Q$ and $c_{\ell}\not=0$. Since $y_{1}-y_{1}\cdot g\in Q$ for all $g\in H$, it follows that $c_{\ell}\in Q^{H}$. Write $\ell=\sum_{i=0}^{n-3}\ell_{i}q^{i}$ with $0\leq\ell_{i}<q$. Define $\beta:=\prod_{j=0}^{n-3}\xi_{j}^{\ell_{j}}$. Then $\operatorname{LT}(\beta)=y_{1}^{\ell}x_{1}^{||\ell||}.$ Since $f$ is strongly compliant, $x_{1}^{||\ell||}$ divides $c_{\ell}$. Define $h:=f-c_{\ell}x_{1}^{-||\ell||}\beta$. Observe that $h\in S^{H}$. Since $\beta$ is strongly compliant, $h$ is strongly compliant with $\deg_{y_{1}}(h)<\ell$. Therefore, using the induction hypothesis, $h\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$. Thus $f=h+c_{\ell}x_{1}^{-||\ell||}\beta\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$. ∎
We will prove by induction on $m$ that $N(y_{1})$ is compliant. As a consequence, $x_{1}N(y_{1})-\xi_{n-2}$ is strongly compliant. Lemma 5.2 then follows from Lemma 5.5.
We first observe that $N(y_{1})$ is compliant for $m=1$. For ${{\text{O}^{+}_{{2}}(\mathbb{F}_{q})}}$, we have $N(y_{1})=y_{1}$, which is trivially compliant.
The following two lemmas are used in the proof of the induction step.
Lemma 5.6.
Suppose that $q-1$ divides $k$, then $q-1$ divides $||k||_{q}$.
Proof.
Write $k=\sum_{s}k_{s}q^{s}$ where $10\leq k_{s}\leq q-$ 1 for all $s$. Since $q^{s}\equiv 1\pmod{(q-1)}$ for all $s$ we have
$0\equiv k=\sum_{s}k_{s}q^{s}\equiv\sum_{s}k_{s}=||k||_{q}\pmod{(q-1)}$
as required. ∎
Define $T:=S[t,u]$ and extend the definition of compliance to $T$. Given $f(y_{1},t)\in T$, we define
$\mathcal{T}_{(y_{1},u,t,x_{1})}(f):=\prod_{a\in\mathbb{F}_{q}}f(y_{1}+au,t-ax_{1})\in T\ .$
Lemma 5.7.
If $f$ is compliant then $\mathcal{T}_{(y_{1},u,t,x_{1})}(f)$ is compliant.
Proof.
We work in the ring $\widetilde{T}:=T[a_{1},a_{2},\dots,a_{q}]$ where $a_{1},a_{2},\dots,a_{q}$ are $q$ new indeterminants. Write $f=\sum_{k}\sum_{j}c_{k,j}y_{1}^{k}t^{j}$. We have $\nu(c_{k,j})\geq||k||-1$ for all $k$ and $j$. Consider
$\widetilde{f}$ $=\prod_{s=1}^{q}f(y_{1}+a_{s}u,t-a_{s}x_{1})\in\widetilde{T}$
$=\prod_{s=1}^{q}\left(\sum_{k}\sum_{j}c_{k,j}(y_{1}+a_{s}u)^{k}(t-a_{s}x_{1})^{j}\right)$
[…]
Thus $\widetilde{f}$ is a linear combination of expressions of the form
[…]
which can be written as
[…]
where $k=\sum_{s=1}^{q}(k_{s}-i_{s})$, $i=\sum_{s=1}^{q}i_{s}$, $j=\sum_{s=1}^{q}(j_{s}-\ell_{s})$ and $\ell=\sum_{s=1}^{q}\ell_{s}$.
Clearly $\widetilde{f}$ is invariant under any permutation of the $a_{s}$. Thus we may express $\widetilde{f}$ using the elementary symmetric functions $E_{i}(a_{1},a_{2},\dots,a_{q})$ of degree $i$ in the $a_{i}$, i.e., $\widetilde{f}=\sum_{\vec{e}}h_{(e_{1},e_{2},\dots,e_{q})}E_{1}^{e_{1}}E_{2}^{e_{2}}\cdots E_{q}^{e_{q}}$ with $h_{(e_{1},e_{2},\dots,e_{q})}\in T$. When we specialize the $a_{i}$ to the elements of the field $\mathbb{F}_{q}=\{\alpha_{1},\alpha_{2},\dots,\alpha_{q}\}$, $\widetilde{f}$ specializes to $\mathcal{T}_{(y_{1},u,t,x_{1})}(f)$.
Note that $E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})=0$ for $i\neq q-1$ and $i\geq 1$. This follows from
$X^{q}-X=\prod_{a\in\mathbb{F}_{q}}(X-a)=\sum_{i=0}^{q}(-1)^{i}E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})X^{q-i}$
Therefore in the expression for $\widetilde{f}$ it suffices to consider only terms of the form $h_{(0,0,\dots,0,e,0)}E_{q-1}^{e}$.
Terms occurring in $h_{(0,0,\dots,0,e,0)}$ satisfy $i+\ell=e(q-1)$. The term $h_{(0,0,\dots,0,0,0)}$ when $i=\ell=0$ is just $h_{(0,0,\dots,0,0,0)}=\prod_{s=1}^{q}f(y_{1},t)=f(y_{1},t)^{q}$. Since $||k^{q}||=||k||$ it is clear that $h_{(0,0,\dots,0,0,0)}$ is compliant.
[…]
For $h_{(0,0,\dots,0,e,0)}$ with $e\geq 1$ we have $i+\ell=e(q-1)$ is a positive multiple of $q-1$. Thus $||i+\ell||$ is also a positive multiple of $q-1$ by Lemma 5.6. In particular $||i+\ell||\geq q-1$.
Thus
$||k||=||\sum_{s=1}^{q}k_{s}-i_{s}||\leq\sum_{s=1}^{q}||k_{s}-i_{s}||=\sum_{s=1}^{q}(||k_{s}||-||i_{s}||)\ .$
Therefore
$\ell+\sum_{s=1}^{q}||k_{s}||$ $\geq\ell+||k||+\sum_{s=1}^{q}||i_{s}||\geq\ell+||k||+||\sum_{s=1}^{q}i_{s}||=\ell+||k||+||i||$
$\geq||\ell||+||k||+||i||\geq||k||+||i+\ell||\geq||k||+(q-1)$
Thus $\nu(u^{i}t^{j}x_{1}^{\ell}\prod_{s=1}^{q}c_{k_{s},j_{s}})\geq\ell+\sum_{s=1}^{q}(||k_{s}||-1)\geq||k||-1$ as required. ∎
Let $\overline{H}$ denote the pointwise stabiliser of $\{f_{m},e_{m}\}$ in $H$ and observe that $\overline{H}$ also stabilises $\{y_{m},x_{m}\}$. It is easy to see that $\overline{H}$ is a normal subgroup of $H$. Furthermore, we can identify $H/\overline{H}$ with the subgroup of $H$ consisting of elements which stabilise all variables except $y_{1}$, $y_{m}$, and $x_{m}$. Observe that
$N(y_{1})=\prod_{a,b\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+by_{m}+ax_{m}-abx_{1})$
where $N^{\overline{H}}(v)$ denotes the orbit product of $v$ over $\overline{H}$. By the induction hypothesis, $N^{\overline{H}}(y_{1})$ is compliant. Since $\deg_{x_{m}}(N^{\overline{H}}(y_{1}))=0$, we can apply Lemma 5.7 to $N^{\overline{H}}(y_{1})$ with $u=t=x_{m}$ to show that
$M(y_{1},x_{m}):=\mathcal{T}_{(y_{1},x_{m},x_{m},x_{1})}(N^{\overline{H}}(y_{1}))=\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+ax_{m})$
is compliant. Applying Lemma 5.7 to $M(y_{1},x_{m})$ with $u=y_{m}$, we see that
$\mathcal{T}_{(y_{1},y_{m},x_{m},x_{1})}(M)$ $=\prod_{b\in\mathbb{F}_{q}}M(y_{1}+by_{m},x_{m}-bx_{1})$
$=\prod_{b\in\mathbb{F}_{q}}\bigg{(}\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+by_{m}+a(x_{m}-bx_{1})\bigg{)}$
$=\prod_{b\in\mathbb{F}_{q}}\prod_{a\in\mathbb{F}_{q}}N^{\overline{H}}(y_{1}+by_{m}+ax_{m}-abx_{1})$
$=N(y_{1})$
is compliant.
Since $N(y_{1})$ is compliant, $x_{1}N(y_{1})$ is strongly compliant. Thus $x_{1}N(y_{1})-\xi_{n-2}$ is strongly compliant. Since $\deg_{y_{1}}(x_{1}N(y_{1})-\xi_{n-2})<q^{n-2}$, using Lemma 5.5, we see that
$x_{1}N(y_{1})-\xi_{n-2}\in Q^{H}[\xi_{i}\mid n=0,\ldots,n-3].$
This completes the proof of Lemma 5.2. Proving Lemma 5.2 completes the proof of Lemma 5.3 which in turn completes the proof of Theorem 5.4.
5.8. Computing the Sylow Invariants
Recall that $H$ is a normal subgroup of $P_{m}$. Further more, identifying $P_{m-1}$ with pointwise stabiliser of $\{x_{1},y_{1}\}$ in $P_{m}$, we have $P_{m}=P_{m-1}H$. Therefore $S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}$. Let $\mathcal{B}_{H}$ denote the monomial factors of $\prod_{i=0}^{n-3}\xi_{i}^{q-1}$. Using the results of Section 5.1, we see that $S_{m}^{H}$ is the free module over $Q^{H}[N(y_{1})]$ with basis given by the elements of $\mathcal{B}_{H}$. Thus
$S_{m}^{H}=\bigoplus_{\gamma\in\mathcal{B}_{H}}Q^{H}[N(y_{1})]\gamma.$
Since the elements of $\mathcal{B}_{H}$ are $P_{m}$-invariant, we have
$S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}=\bigoplus_{\gamma\in\mathcal{B}_{H}}(Q^{H}[N(y_{1})])^{P_{m-1}}\gamma.$
Write $\overline{S}$ for the subalgebra of $Q^{H}$ constructed by omitting the generator $x_{1}$. Then $Q^{H}[N(y_{1})]=\overline{S}\otimes\mathbb{F}_{q}[x_{1},N(y_{1})]$ and
$(Q^{H}[N(y_{1})])^{P_{m-1}}=\overline{S}^{P_{m-1}}\otimes\mathbb{F}_{q}[x_{1},N(y_{1})].$
The action of $P_{m-1}$ on $\overline{S}$ is equivalent to the action of $P_{m-1}$ on $S_{m-1}$. This gives an algebra isomorphism from $S_{m-1}^{P_{m-1}}$ to $\overline{S}^{P_{m-1}}$ which takes elements of degree $d$ to elements of degree $qd$. This isomorphism is induced by $\varphi:=\psi_{1}\circ\sigma$, where $\sigma$ is the algebra monomorphism from $S_{m-1}$ to $S_{m}$ which takes $y_{i}$ to $y_{i+1}$, $x_{i}$ to $x_{i+1}$. Note that the isomorphism takes $\xi_{i}\in S_{m-1}$ to $\psi_{1}(\xi_{i})\in S_{m}$.
Theorem 5.9.
$S_{m}^{P_{m}}$ is Cohen-Macaulay and is generated by the orbit products of the variables and the $\xi_{i}$.
Proof.
The proof is by induction on $m$. For $m=2$, we have $P_{m}=H$ and the result follows from Section 5.1.
For the induction step, first observe that $S_{m}^{P_{m}}$ is generated by $x_{1}$, $N(y_{1})$, elements of $\mathcal{B}_{H}$ and generators for $\overline{S}^{P_{m-1}}$. The isomorphism from $S_{m-1}^{P_{m-1}}$ to $\overline{S}^{P_{m-1}}$ takes orbit products to orbit products and $\xi_{i}$ to $\psi_{1}(\xi_{i})\in\mathbb{F}_{q}[x_{1},\xi_{j}\mid j\leq i+1]$. Thus $S_{m}^{P_{m}}$ is generated by orbit products and the $\xi_{i}$. By induction, $S_{m-1}^{P_{m-1}}$ is a free module over the algebra generated by the orbit products of the variables. Let $\mathcal{C}$ denote the image of the module generators in $\overline{S}^{P_{m-1}}$. Then $\mathcal{B}_{H}\mathcal{C}$ is the set of module generators for $S_{m}^{P_{m}}$ over the algebra generated by the orbit products of the variables and $S_{m}^{P_{m}}$ is Cohen-Macaulay. ∎
Define algebra maps $\psi_{j}:S_{m}\to S_{m}$ by
$\psi_{j}(v):=\prod_{u\in U_{j}}(v-u)$
for $v$ homogeneous of degree one, where $U_{j}$ is the span of $\{x_{1},\ldots,x_{j}\}$. These maps $\psi_{j}$ were introduced and used effectively by Ferreira and Fleischmann in [14].
Lemma 5.10.
In the lexicographic term order $\operatorname{LT}(\psi_{j}(\xi_{i}))=y_{j+1}^{q^{i+j}}x_{j+1}^{q^{j}}$.
Proof.
The linear forms $x_{1},x_{2},\dots,x_{j}$ lie in the kernel of $\psi_{j}$. Furthermore for $k>j$ and $a,b$ positive integers $\operatorname{LT}(\psi_{j}(y_{k}^{a}x_{k}^{b}))=y_{k}^{q^{j}a}x_{k}^{q^{j}b}$. From these observations the result follows. ∎
Lemma 5.11.
$\varphi\circ\psi_{j}=\psi_{j+1}\circ\sigma$ and $\varphi(\psi_{j}(\xi_{i}))=\psi_{j+1}(\xi_{i})$.
Proof.
For $j=0$, the first equation is the definition of $\varphi$ and the second follows from the observation that $x_{1}$ is in the kernel of $\psi_{1}$.
Suppose $j>0$. Both $\varphi\circ\psi_{j}$ and $\psi_{j+1}\circ\sigma$ are algebra maps so to prove the first equation it is sufficient to show that $\varphi(\psi_{j}(v))=\psi_{j+1}(\sigma(v))$ when $\deg(v)=1$. We have
$\varphi(\psi_{j}(v))=\psi_{1}\left(\sigma\left(\prod_{u\in U_{j}}(v-u)\right)\right)=\psi_{1}\left(\prod_{u\in U_{j}}(\sigma(v)-\sigma(u))\right).$
However, $U_{j+1}=U_{1}\oplus\sigma(U_{j})$. Therefore
$\psi_{1}\left(\prod_{u\in U_{j}}(\sigma(v)-\sigma(u))\right)=\prod_{u\in U_{j+1}}(\sigma(v)-u)=\psi_{j+1}(\sigma(v))$
as required. For the second equation, we have $\varphi(\psi_{j}(\xi_{i}))=\psi_{j+1}(\sigma(\xi_{i}))=\psi_{j+1}(\xi_{i})$, since $x_{1}$ is in the kernel of $\psi_{j+1}$. ∎
Example 5.12.
In this example we consider the Sylow invariants for ${{\text{O}^{+}_{{6}}(\mathbb{F}_{q})}}$. It follows from Theorem 5.9 that $S_{3}^{P_{3}}$ is generated by the homogeneous system of parameters $\mathcal{H}_{0}:=\{N(y_{1}),N(y_{2}),N(y_{3}),N(x_{3}),N(x_{2}),x_{1}\}$ and $\{\xi_{0},\xi_{1},\xi_{2},\xi_{3}\}$. Using the lexicographic order the lead monomials are
$y_{1}^{q^{4}},y_{2}^{q^{3}},y_{3}^{q^{2}},x_{3}^{q^{2}},x_{2}^{q},x_{1},y_{1}x_{1},y_{1}^{q}x_{1},y_{1}^{q^{2}}x_{1},y_{1}^{q^{3}}x_{1}.$
The initial tête-a-têtes are
$\xi_{0}^{q}-x_{1}^{q-1}\xi_{1},\,\xi_{1}^{q}-x_{1}^{q-1}\xi_{2},\,\xi_{2}^{q}-x_{1}^{q-1}\xi_{3},\,\xi_{3}^{q}-x_{1}^{q}N(y_{1}).$
Subducting $\xi_{0}^{q}-x_{1}^{q-1}\xi_{1}$ gives $\psi_{1}(\xi_{0})$, which has lead monomial $y_{2}^{q}x_{2}^{q}$. Subducting $\xi_{1}^{q}-x_{1}^{q-1}\xi_{2}$ gives $\psi_{1}(\xi_{1})$, which has lead monomial $y_{2}^{q^{2}}x_{2}^{q}$. This suggests that a Khovanskii bases for $S_{3}^{P_{3}}$ should include $\psi_{1}(\xi_{0})$ and $\psi_{1}(\xi_{1})$. We know that $S_{2}^{P_{2}}$ is generated by $\xi_{0}$, $\xi_{1}$ and the orbit products of the variables. We also know that this set is a Khovanskii basis. Using the isomorphism from $S_{2}^{P_{2}}$ to $\overline{S}^{P_{2}}$, we conclude that $\{N(y_{2}),N(y_{3}),N(x_{3}),N(x_{2}),\psi_{1}(\xi_{0}),\psi_{1}(\xi_{1})\}$ is a Khovanskii basis for the subalgebra $\overline{S}^{p_{2}}\subset S_{3}^{P_{3}}$. Therefore the tête-a-têtes $\psi_{1}(\xi_{0})^{q}-N(x_{2})^{q-1}\psi_{1}(\xi_{1})$, $\psi_{1}(\xi_{1})^{q}-N(x_{2})^{q}N(y_{2})$ and $\psi_{1}(\xi_{2})-N(x_{2})N(y_{2})^{q}$ subduct to zero. As a result, the tête-a-têtes $(\xi_{0}^{q}-x_{1}^{q-1}\xi_{1})^{q}$, $(\xi_{1}^{q}-x_{1}^{q-1}\xi_{2})^{q}$and $\xi_{2}^{q}-x_{1}^{q-1}\xi_{3}$ subduct to zero. To prove that adjoining $\psi_{1}(\xi_{0})$ and $\psi_{1}(\xi_{1})$ to our generating set gives a Khovanskii basis, we need only show that the tête-a-tête $\xi_{3}^{q}-x_{1}^{q}N(y_{1})$ subducts to zero. We know that this tête-a-tête subducts to zero over the Hook invariants but it is not immediately clear that this subduction can be constructed over the Sylow invariants. It follows from the proof of Theorem 5.9 that $S_{3}^{P_{3}}$ is the free module over the algebra generated by $\mathcal{H}_{0}$ with module generators
$\mathcal{B}_{H}\mathcal{C}=\{\xi_{0}^{a_{0}}\xi_{1}^{a_{1}}\xi_{2}^{a_{2}}\xi_{3}^{a_{3}}\psi_{1}(\xi_{0})^{b_{1}}\psi_{1}(\xi_{1})^{b_{2}}\mid a_{i}<q,\,b_{i}<q\}.$
The lead monomials are given by
$\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})=\{x_{1}^{a_{0}+a_{1}+a_{2}+a_{3}}y_{1}^{a_{0}+a_{1}q+a_{2}q^{2}+a_{3}q^{3}}x_{2}^{q(b_{1}+b_{2})}y_{2}^{qb_{1}+q^{2}b_{2}}\mid a_{i}<q,\,b_{i}<q\}.$
Therefore $\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})$ generates a free module of rank $q^{6}$ over the algebra generated by $\operatorname{LT}(\mathcal{H}_{0})$. This module is a subalgebra of $\operatorname{LT}(S_{3}^{P_{3}})$. By comparing Hilbert series and using the fact that $\operatorname{LT}(S_{3}^{P_{3}})$ and $S_{3}^{P_{3}}$ have the same Hilbert series, we see that $\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S_{3}^{P_{3}}$. As a consequence, $S_{3}^{P_{3}}$ is a complete intersection with relations constructed by subducting the tête-a-têtes $\xi_{0}^{q^{2}}-(x_{1}^{q-1}\xi_{1})^{q}$, $\xi_{1}^{q^{2}}-(x_{1}^{q-1}\xi_{2})^{q}$, $\xi_{2}^{q}-x_{1}^{q-1}\xi_{3}$ and $\xi_{3}^{q}-x_{1}^{q}N(y_{1})$. Furthermore, we could replace the module generators $\mathcal{B}_{H}\mathcal{C}$ with $\{\xi_{0}^{e_{0}}\xi_{1}^{e_{1}}\xi_{2}^{e_{2}}\xi_{3}^{e_{3}}\mid e_{0},e_{1}<q^{2},\,e_{2},e_{3}<q\}$.
Let $\mathcal{H}_{0}$ denote the system of parameters given by the orbit products of the variables and define $\mathcal{B}_{H}$ and $\mathcal{C}$ as in the proof of Theorem 5.9.
Theorem 5.13.
Using the lexicographic order, $\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S_{m}^{P_{m}}$. Furthermore $S_{m}^{P_{m}}$ is a complete intersection.
Proof.
Applying Lemma 5.11 inductively shows that the elements of $\mathcal{B}_{H}\mathcal{C}$ are of the form $\prod\psi_{j}(\xi_{i})^{e_{ij}}$ with $e_{ij}<q$. Using Lemma 5.10, careful book-keeping shows that $\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})$ has the same cardinality as $\mathcal{B}_{H}\mathcal{C}$ and that $\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})$ generates a free module over the algebra generated by $\operatorname{LT}(\mathcal{H}_{0})$. This module is a subalgebra of $\operatorname{LT}(S_{m}^{P_{m}})$. By comparing Hilbert series and using the fact that $\operatorname{LT}(S_{m}^{P_{m}})$ and $S_{m}^{P_{m}}$ have the same Hilbert series, we see that $\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S_{m}^{P_{m}}$. Let $\widetilde{\mathcal{B}_{H}}$ denote the subset of $\mathcal{B}_{H}\mathcal{C}$ consisting of elements of the form $\psi_{j}(\xi_{i})$. Since the algebra generated by $\operatorname{LT}(\mathcal{H}_{0}\cup\widetilde{\mathcal{B}_{H}})$ coincides with the algebra generated by $\operatorname{LT}(\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C})$, we see that $\mathcal{H}_{0}\cup\widetilde{\mathcal{B}_{H}}$ is also a Khovanskii basis for $S_{m}^{P_{m}}$.
Note that ~$\widetilde{\mathcal{B}}_{H}=\{\psi_{j}(\xi_{i})\mid 0\leq i\leq n-3-2j\}$. The non-trivial tête-a-têtes are $\psi_{j}(\xi_{i})^{q}-\psi_{j}(\xi_{i+1})N(x_{j+1})^{q-1}$ for $i<n-3-2j$ and $\psi_{j}(\xi_{i})^{q}-N(y_{j+1})N(x_{j+1})^{q}$ for $i=n-3-2j$. Since $\mathcal{H}_{0}\cup\widetilde{\mathcal{B}_{H}}$ is a Khovanskii basis, each of these tête-a-têtes subducts to zero. The resulting relations are independent and generate the ideal of relations. The number of generators minus the number of independent relations is $n$, proving that $S_{m}^{P_{m}}$ is a complete intersection. ∎
Remark 5.14.
It follows from the proof of Theorem 5.13 that $\mathcal{B}_{H}\mathcal{C}$ consists of the monomial factors of $\prod\{\psi_{j}(\xi_{i})^{q-1}\mid 0\leq i\leq n-3-2j\}$.
Corollary 5.15.
The monomial factors of
$\Gamma_{0}=\prod_{i=0}^{n-3}\xi_{i}^{e_{i}}\text{ where }e_{i}=q^{m-1-\lfloor i/2\rfloor}-1\ .$
form a basis for $S_{m}^{P_{m}}$ as a free module over $\mathbb{F}_{q}[\mathcal{H}_{0}]$.
Proof.
Using Theorem 5.13 and its proof, each of the tête-a-têtes $\xi_{n-3}^{q}-x_{1}^{q}N(y_{1})$, $\xi_{n-4}^{q}-x_{1}^{q-1}\xi_{n-3}$, $\xi_{n-5}^{q^{2}}-x_{1}^{q-1}\xi_{n-4}^{q}$, $\xi_{n-6}^{q^{2}}-x_{1}^{q-1}\xi_{n-5}^{q},\dots,\xi_{1}^{q^{m-1}}-x_{1}^{q-1}\xi_{2}^{q^{m-2}}$, and $\xi_{0}^{q^{m-1}}-x_{1}^{q-1}\xi_{1}^{q^{m-2}}$ subducts to zero. Applying Theorem 5.9 and comparing the number of monomial factors of $\Gamma_{0}$ with the number of module generators for $S_{m}^{P_{m}}$ over $\mathbb{F}_{q}[\mathcal{H}_{0}]$ yields the result. ∎
Remark 5.16.
The rank of $S_{m}^{P_{m}}$ as a free $\mathbb{F}_{q}[\mathcal{H}_{0}]$-module is $q^{m(m-1)}$ which surprisingly is also $|P_{m}|$.
Corollary 5.17.
$S_{m}^{P_{m}}$ is minimally generated by $\mathcal{H}_{0}\cup\{\xi_{0},\xi_{1},\dots,\xi_{n-3}\}$.
Proof.
In [14] it is shown that
$\mathcal{F}(S_{m}^{P_{m}})=\mathbb{F}_{q}(N(y_{m}),N(x_{1}),N(x_{2}),\dots,N(x_{m}),\xi_{0},\xi_{1},\dots,\xi_{m-2})$
and hence this set of invariants is algebraically independent. Clearly
$\mathcal{V}(N(y_{m}),N(x_{1}),N(x_{2}),\dots,N(x_{m}),\xi_{0},\xi_{1},\dots,\xi_{m-2})=\operatorname{Span}_{\overline{\mathbb{F}_{q}}}\{e_{1},e_{2},\dots,e_{m-1}\}.$
The remaining $\xi_{i}$ all vanish on this set and so to cut this variety down to a point (the origin) we need to add in the remaining $m-1$ norms. It remains to show that $\xi_{m-1},\xi_{m},\dots,\xi_{n-3}$ are also necessary as generators.
Otherwise, one of these $\xi_{i}$ can be expressed as a polynomial $f$ in the other $32n-$ 3 generators. Since $\operatorname{LT}(\xi_{i})=y_{1}^{q^{i}}x_{1}$ this must also be the lead term of $f$ as an element of $S_{m}$. This lead term cannot be expressed as a product of the lead terms of the other $32m-$ 3 generators and thus it must arise from a tête-a-tête. This tête-a-tête must be in degree $q^{i}+1$ and involve at least one generator whose lead term is greater than $y_{1}^{q^{i}}x_{1}$. But any of these $32m-$ 3 generators whose lead term is greater than $y_{1}^{q^{i}}x_{1}$ also has degree greater than $q^{i}+1$. Hence no such tête-a-tête can exist. ∎
5.18. Computing the Borel Invariants
Let $B$ denote the Borel subgroup of upper triangular matrices. The Sylow subgroup $P$ is a normal subgroup of $B$ and we identify $B/P$ with the subgroup of diagonal matrices. Let $\mathcal{H}_{0}$ denote the homogeneous system of parameters for $S^{P}$ given by the orbit products over $P$ of the variables and let $\mathbb{F}_{q}[\mathcal{H}_{0}]$ denote the algebra generated by $\mathcal{H}_{0}$. Define $\mathcal{B}_{H}$ and $\mathcal{C}$ as in the proof of Theorem 5.9. Then
$S^{P}=\bigoplus_{\gamma\in\mathcal{B}_{H}\mathcal{C}}\mathbb{F}_{q}[\mathcal{H}_{0}]\gamma.$
Lemma 5.19.
Suppose $w\in S$ with $w$ homogeneous of degree one and $b\in B$ with $w\cdot b=cw$ for some scalar $c$. If the $P$-orbit product $N(w)$ has degree $q^{\ell}$ then $N(w)\cdot b=c^{q^{\ell}}N(w)$.
Proof.
Since $P$ is normal in $B$, for every $g\in P$, we have $\overline{g}:=b^{-1}gb\in P$. Thus
$N(w)b$ =($=\left(\prod\{wg\mid g\in P\}\right)b=\prod\{wgb\mid g\in P\}=\prod\{wb\overline{g}\mid g\in P\}$
=$=\prod\{cw\overline{g}\mid g\in P\}=c^{q^{\ell}}\prod\{wg\mid g\in P\}=c^{q^{\ell}}N(w)$
as required. ∎
Since the elements of $\mathcal{H}_{0}$ are orbit products over $P$, it follows from Lemma 5.19 that $B/P$ acts on $\mathbb{F}_{q}[\mathcal{H}_{0}]$. Arguing as in the proof of Theorem 5.13, the elements of $\mathcal{B}_{H}\mathcal{C}$ are of the form $\psi_{j}(\xi_{i})$. Since $\psi_{j}(\xi_{i})$ is Borel invariant, we have
$S^{B}=(S^{P})^{B/P}=\bigoplus_{\gamma\in\mathcal{B}_{H}\mathcal{C}}\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}\gamma.$
Therefore $S^{B}$ is a free module over $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$ and to compute $S^{B}$, we need only compute $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$.
Theorem 5.20.
$S^{B}$ is a complete intersection. Furthermore $\widetilde{\mathcal{H}_{0}}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S^{B}$, using the lexicographic order, where
$\widetilde{\mathcal{H}_{0}}:=\{N(y_{i})^{q-1},N(x_{i})^{q-1},N(y_{i})N(x_{i})\mid i=1,\ldots m\}.$
Proof.
The proof follows from the computation of $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$. It follows from Lemma 5.19 that $N(y_{i})^{q-1}$ and $N(x_{i})^{q-1}$ are Borel invariant. Furthermore, $B$ contains the linear transformation which scales $y_{i}$ by $c\in\mathbb{F}_{q}$, takes $x_{i}$ to $c^{-1}x_{i}$, and fixes the other variables. Therefore no smaller power of $N(y_{i})$ or $N(x_{i})$ is Borel invariant. Any element of $B/P$ which scales $y_{i}$ by $c$ must scale $x_{i}$ by $c^{-1}$. Therefore $N(y_{i})N(x_{i})$ is Borel invariant. It follows from this that $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$ is the complete intersection generated by $\widetilde{\mathcal{H}_{0}}$ subject to the relations $(N(y_{i})N(x_{i}))^{q-1}=N(y_{i})^{q-1}N(x_{i})^{q-1}$ and $\widetilde{\mathcal{H}_{0}}$ and is a Khovanskii basis for $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$. Since there are no new tête-a-têtes involving $\mathcal{B}_{H}\mathcal{C}$, $\widetilde{\mathcal{H}_{0}}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S^{B}$, as required. ∎
The Weyl group, $W$, of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ is isomorphic to the wreath product $C_{2}\wr\Sigma_{m}$. We can identify $W$ with a subgroup of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$: $\Sigma_{m}$ acts by permuting the hyperbolic pairs and the additional generator, say $\omega$, interchanges $y_{m}$ and $x_{m}$ while fixing the other variables.
Let $\mathcal{R}$ denote the Reynolds operator from $S^{B}$ to $S^{G}$. Since the order of $B$ is $q^{m(m-1)}(q-1)^{m}$, the index
$[G:B]=2\frac{q^{m}-1}{q-1}\prod_{j=1}^{m-1}\frac{q^{2j}-1}{q-1}\equiv_{(p)}2$
and
$\mathcal{R}(f)=\frac{1}{2}\operatorname{Tr}_{B}^{G}(f)=\frac{1}{2}\sum_{Bg\in B\backslash G}fBg$
To compute the Reynolds operator we need to choose left coset representatives for $B$ in ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. We use the double coset partition to do this. The double cosets, $BwB$ for $w\in W$, partition ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ so we can choose coset representatives of the form $wb$ for $w\in W$ and $b\in B$. Observe that the number of left cosets in the double coset $BwB$ is given by the index $[B:B\cap w^{-1}Bw]=[P:P\cap w^{-1}Pw]$, which divides $q^{m(m-1)}$.
Theorem 5.21.
Using the lexicographic order on $S_{m}$,
$\operatorname{LT}\left(\mathcal{R}(N(y_{1})^{q-1})\right)=y_{1}^{(q-1)q^{n-2}}.$
Proof.
Note that $y_{1}^{(q-1)q^{n-2}}$ is the largest monomial of degree $(q-1)q^{n-2}$. The elements of $W$ permute the variables. Every term of $N(y_{1})$ is divisible by $y_{1}$. Therefore the lead term of $N(y_{1})^{q-1}w$ is $y_{1}^{(q-1)q^{n-2}}$ if and only if $w$ stablises $y_{1}$. Since elements of $B$ do not increase the $y_{1}$-degree, the only coset representatives that contribute to the coefficient of $y_{1}^{(q-1)q^{n-2}}$ in $\mathcal{R}(N(y_{1})^{q-1})$ are those of the form $wb$ were $w$ stablises $y_{1}$. If $w$ stabilises $y_{1}$, then all of the coset representatives of the form $wb$ make the same contribution to the coefficient. The number of coset representatives associated to a double coset is either $1$, if $P\cap w^{-1}Pw=P$, or congruent to 00 modulo $p$. The only elements of $W$ for which $P\cap w^{-1}Pw=P$ are $1$ and $\omega$. Therefore the coefficient of $y_{1}^{(q-1)q^{n-2}}$ in $\mathcal{R}(N(y_{1})^{q-1})$ is $(1/2)2=1$. ∎
6. The Invariants of ${{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}$
The ring of invariants for ${{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}$ was computed by Huah Chu [8]. In this section we provide an alternative computation and show that Theorem 4.6 is satisfied for $m=2$. This provides the base case for the inductive proof of the theorem.
The group ${{\text{O}^{+}_{{2}}(\mathbb{F}_{q})}}$ has order $2(q-1)$ and the ring of invariants is generated by $\xi_{0}=y_{1}x_{1}=u_{1}$ and $d_{1,1}=y_{1}^{q-1}+x_{1}^{q-1}=\xi_{1}/\xi_{0}$.
Lemma 6.1.
$u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}\in R_{1}$. Furthermore
$c_{2,2}=\xi_{0}\xi_{1}^{q-1}+\xi_{0}^{q+2}\xi_{1}^{q-3}-\sum_{j=2}^{(q-1)/2}(-1)^{j}\xi_{1}^{q-1-2j}\xi_{0}^{1+j(q+1)}\frac{(q-2-j)!}{j!((q-1)-2j)!}$
and $\ker(\Phi_{2,1})=u_{2}R_{2}$.
Proof.
Working in $S_{1}$, we have $\xi_{1}=\xi_{0}d_{1,1}$ where $d_{1,1}=y^{q-1}+x^{q-1}$. Applying $\mathcal{P}^{q}$, using the Cartan identity and the stability condition, gives $\xi_{2}=\xi_{1}\mathcal{P}^{q-1}(d_{1,1})+\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})=\xi_{1}d_{1,1}^{q}+\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})$. Multiplying by $\xi_{0}^{q}$ gives $\xi_{0}^{q}\xi_{2}=\xi_{1}^{q+1}+\xi_{0}^{2q}\mathcal{P}^{q-2}(d_{1,1})$. We will show that $\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})\in R_{1}$. From this it, follows that the relation lifts to $S_{2}$, giving a non-zero element of $\ker(\Phi_{2,1})$. Using Lemma 4.2, the element is divisible by $u_{2}$ Comparing lead terms using the lexicographic order in $S_{2}$ gives $u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}=-\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})$. A simple calculation gives $\mathcal{P}^{q-2}(d_{1,1})=-(y^{q^{2}-2q+1}+x^{q^{2}-2q+1})$. Taking $a=y^{q-1}$ and $b=x^{q-1}$, we have $-\mathcal{P}^{q-2}(d_{1,1})=a^{q-1}+b^{q-1}$. We can write the symmetric polynomial $a^{q-1}+b^{q-1}$ as a polynomial in $\sigma_{1}=a+b$ and $\sigma_{2}=ab$. A closed form for this expression over the integers is
$a^{q-1}+b^{q-1}=(q-1)\sum_{j=0}^{(q-1)/2}(-1)^{j}\sigma_{1}^{q-1-2j}\sigma_{2}^{j}\frac{(q-2-j)!}{j!((q-1)-2j)!},$
see [16] formula (1) or (10) (note there is a typographical error in [16](10) where $i+j+1$ is written but $i+j-1$ is intended). Observe that $\sigma_{2}=\xi_{0}^{q-1}$, $\sigma_{1}=d_{1,1}$ and $\xi_{0}\sigma_{1}=\xi_{1}$. Multiplying the above expression by $\xi_{0}^{q}$ and substituting gives
$\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})=\sum_{j=0}^{(q-1)/2}(-1)^{j}\xi_{1}^{q-1-2j}\xi_{0}^{1+j(q+1)}\frac{(q-2-j)!}{j!((q-1)-2j)!}\in R_{1},$
as required. ∎
Lemma 6.2.
$(-1)^{j}(q-1)\frac{(q-2-j)!}{j!((q-1)-2j)!}\equiv_{(p)}\operatorname{{Cat}}(j)$
where $\operatorname{{Cat}}(j)=\frac{1}{j+1}\binom{2j}{j}$ is the $j^{\text{th}}$ Catalan number. In particular
$c_{2,2}=\sum_{j=0}^{(q-1)/2}\operatorname{{Cat}}(j)\,\xi_{0}^{j(q+1)+1}\xi_{1}^{q-1-2j}.$
Proof.
Both of the expressions $(-1)^{j}(q-1)\frac{(q-2-j)!}{j!((q-1)-2j)!}$ and $\frac{1}{j+1}\binom{2j}{j}$ are integers. We show they are equal when viewed as elements of $\mathbb{F}_{p}$:
$(-1)^{j}(q-1)\frac{(q-2-j)!}{j!((q-1)-2j)!}$ $=(-1)^{j+1}\frac{(q-2-j)!}{j!((q-1)-2j)!}$
$=\frac{(-1)^{j+1}}{j!}(q-2j)(q-2j+1)\cdots(q-2-j)$
$=\frac{(-1)^{j+1}}{j!}(-2j)(-2j+1)\cdots(-2-j)$
$=\frac{(-1)^{j+1}}{j!}(-1)^{j-1}(2j)(2j-1)\cdots(j+2)$
$=\frac{(2j)!}{j!(j+1)!}=\frac{1}{j+1}\binom{2j}{j}\ .\qed$
Note that $u_{2}=(\xi_{2}+c_{2,2})u_{1}^{q}-\xi_{1}(u_{1}d_{1,1})^{q}$. Therefore the displayed equation in part(a) of Theorem 4.6 holds for $m=2$.
Lemma 6.3.
(a) $\mathcal{P}^{1}(c_{2,2})=\xi_{1}^{q}$.
(b) $\mathcal{P}^{i}(c_{2,2})=0$ for $1<i<q$.
(c) $\mathcal{P}^{q^{2}}(c_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}$.
(d) $\mathcal{P}^{q^{2}-q}(c_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{0}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2}$.
Proof.
Since $\mathcal{P}^{1}(u_{2})=0$, we have
$0=\mathcal{P}^{1}(\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1})=\xi_{0}^{q}(\xi_{1}^{q}+\mathcal{P}^{1}(c_{2,2}))-2\xi_{0}^{q}\xi_{1}^{q}.$
Therefore $\mathcal{P}^{1}(c_{2,2})=\xi_{1}^{q}$, proving (a). For $1<i<q$, we have
$0=\mathcal{P}^{i}(\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1})=\xi_{0}^{q}\mathcal{P}^{i}(c_{2,2}),$
giving $\mathcal{P}^{i}(c_{2,2})=0$. Since $c_{2,2}\equiv_{\langle\xi_{0}^{q}\rangle}\xi_{0}\xi_{1}^{q-1}$, it follows from Lemma 3.6 that
$\mathcal{P}^{q^{2}}(c_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\mathcal{P}^{q^{2}}(\xi_{0}\xi_{1}^{q-1})=\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2},$
proving (c). To prove (d), we first compute
$\mathcal{P}^{q^{2}-q}(\xi_{0}\xi_{1}^{q-1})=\xi_{0}\mathcal{P}^{q^{2}-q}(\xi_{1}^{q-1})+\xi_{1}\mathcal{P}^{q^{2}-q-1}(\xi_{1}^{q-1})+\xi_{0}^{q}\mathcal{P}^{q^{2}-q-2}(\xi_{1}^{q-1}).$
Note that $\mathcal{P}^{q^{2}-q}(\xi_{1}^{q-1})=\xi_{2}^{q-1}$, $\mathcal{P}^{q^{2}-q-1}(\xi_{1}^{q-1})=-2\xi_{0}^{q}\xi_{1}^{q^{2}-2q}$ and
$\mathcal{P}^{q^{2}-q-2}(\xi_{1}^{q-1})=-\xi_{1}^{q^{2}-2q+1}-\xi_{2}\mathcal{P}^{q^{2}-2q-2}(\xi_{1}^{q-2})=-\xi_{1}^{q^{2}-q+q}+4\xi_{2}\xi_{0}^{q}\xi_{1}^{q^{2}-3q}.$
Therefore $\mathcal{P}^{q^{2}-q}(\xi_{0}\xi_{1}^{q-1})=\xi_{0}\xi_{2}^{q-1}-3\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+4\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2}$. Observe that $\mathcal{P}^{q^{2}-q}(\xi_{0}^{q+2}\xi_{1}^{q-3})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\mathcal{P}^{q^{2}-2q}(\xi_{0}^{2}\xi_{1}^{q-3})$ and
$\mathcal{P}^{q^{2}-2q}(\xi_{0}^{2}\xi_{1}^{q-3})=2\xi_{0}^{q}\xi_{1}^{q^{2}-3q+1}-3\xi_{0}^{2q}\xi_{1}^{q^{2}-4q}\xi_{2}.$
Therefore $\mathcal{P}^{q^{2}-q}(\xi_{0}^{q+2}\xi_{1}^{q-3})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}2\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}-3\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}$. Suppose $b$ is homogeneous of degree $q^{2}-4q+1$. Then $\mathcal{P}^{q^{2}-q}(\xi_{0}^{2q}b)\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\mathcal{P}^{q^{2}-2q}(\xi_{0}^{q}b)$ and, using Lemma 3.6, $\mathcal{P}^{q^{2}-2q}(\xi_{0}^{q}b)=\xi_{0}^{q^{2}}\mathcal{P}^{q^{2}-4q}(b)$. Therefore $\mathcal{P}^{q^{2}-q}(\xi_{0}^{2q}b)\equiv_{\langle\xi_{0}^{q^{2}}\rangle}0$ and
$\mathcal{P}^{q^{2}-q}(c_{2,2})$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\mathcal{P}^{q^{2}-q}(\xi_{0}\xi_{1}^{q-1})+\mathcal{P}^{q^{2}-q}(\xi_{0}^{q+2}\xi_{1}^{q-3})$
$\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{0}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2}$
as required. ∎
By Definition, we have $\mathcal{P}^{q^{2}}(u_{2})=u_{2}d_{1,2}$ and $\mathcal{P}^{q^{2}+q}(u_{2})=u_{2}d_{2,2}$.
Lemma 6.4.
(a) $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$ and
$u_{2}d_{1,2}\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{0}^{q}\xi_{3}-\xi_{1}\xi_{2}^{q}+\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1}.$
(b) $u_{2}d_{2,2}-\xi_{1}^{q}\xi_{3}\in R_{2}$ and
$u_{2}d_{2,2}\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{3}-\xi_{2}^{q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-q}\xi_{2}.$
Proof.
Using Lemma 6.1, Lemma 3.5 and the Cartan identity,
$\mathcal{P}^{q^{2}}(u_{2})=\xi_{0}^{q}(\xi_{3}+\mathcal{P}^{q^{2}}(c_{2,2}))-\xi_{1}\xi_{2}^{q}+\xi_{1}^{q}\mathcal{P}^{q^{2}-q}(c_{2,2})+\xi_{0}^{q^{2}}\mathcal{P}^{q^{2}-2q}(c_{2,2}).$
Since $c_{2,2}\in R_{1}$, we have $\mathcal{P}^{i}(c_{2,2})\in R_{2}$ and $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$. Using Lemma 6.3, we have
$u_{2}d_{1,2}=\mathcal{P}^{q^{2}}(u_{2})$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{0}^{q}(\xi_{3}+\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2})-\xi_{1}\xi_{2}^{q}$
$+\xi_{1}^{q}(\xi_{0}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2})$
$\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{0}^{q}\xi_{3}-\xi_{1}\xi_{2}^{q}+\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1}$
as required. Using the stability condition
$\mathcal{P}^{q^{2}+q}(u_{2})=\xi_{1}^{q}(\xi_{3}+\mathcal{P}^{q^{2}}(c_{2,2}))-\xi_{2}^{q+1}-\xi_{1}^{q^{2}+1}+\xi_{0}^{q^{2}}\mathcal{P}^{q^{2}-q}(c_{2,2}).$
Using Lemma 6.3 gives
$\mathcal{P}^{q^{2}+q}(u_{2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{3}-\xi_{2}^{q+1}-\xi_{1}^{q^{2}+1}+\xi_{1}^{q}(\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}).$
Therefore $u_{2}d_{2,2}\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{3}-\xi_{2}^{q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-q}\xi_{2}$, as required. ∎
Lemma 6.5.
(a) $\mathcal{P}^{i}(u_{2})=0$ for $0<i<q$.
(b) $\mathcal{P}^{i}(u_{2})/u_{2}\in R_{2}$ for $i\not\in\{q^{2},q^{2}+q,q^{2}+2q\}$.
Proof.
As an element of $S_{2}$, $u_{2}$ is a product of linear forms. Therefore $u_{2}$ divides $\mathcal{P}^{i}(u_{2})$ in $S_{2}$. Recall $u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}\in R_{1}$. For $0<i<q$, we have $\mathcal{P}^{i}(u_{2})\in R_{1}$. Since $u_{2}$ is linear in $\xi_{2}$, we have $\mathcal{P}^{i}(u_{2})=0$, proving (a). For $i\not\in\{q^{2},q^{2}+q,q^{2}+2q\}$, we have $\mathcal{P}^{i}(u_{2})\in R_{2}$. Therefore $\mathcal{P}^{i}(u_{2})\in\ker(\Phi_{2,1})=u_{2}R_{2}$, proving (b). ∎
Lemma 6.6.
(a) For $0<i<q$, $\mathcal{P}^{i}(d_{1,2})$ and $\mathcal{P}^{i}(d_{2,2})$ lie in $R_{2}$.
(b) For $0<i<q^{3}$, write $i=\ell q^{2}+r$ with $r<q^{2}$. Then $u_{2}^{\ell+1}\mathcal{P}^{i}(d_{1,2})$ and $u_{2}^{\ell+1}\mathcal{P}^{i}(d_{2,2})$ lie in $R_{3}$ with $\xi_{3}$-degree at most $\ell+1$.
Proof.
Suppose $0<i<q$. Using Lemma 6.5, $\mathcal{P}^{i}(u_{2})=0$. Therefore $\mathcal{P}^{i}(u_{2}d_{1,2})=u_{2}\mathcal{P}^{i}(d_{1,2})$ and $\mathcal{P}^{i}(u_{2}d_{2,2})=u_{2}\mathcal{P}^{i}(d_{2,2})$. Since $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$ and $i<q$, we have $\mathcal{P}^{i}(u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3})\in R_{2}$. Since $\mathcal{P}^{1}(\xi_{0}^{q}\xi_{3})=\xi_{0}^{q}\xi_{2}^{q}$ and $\mathcal{P}^{i}(\xi_{0}^{q}\xi_{3})=0$ for $1<i<q$, we see that $\mathcal{P}^{i}(u_{2}d_{1,2})\in R_{2}$. Thus $u_{2}\mathcal{P}^{i}(d_{1,2})\in\ker(\Phi_{2,1})=u_{2}R_{2}$ and $\mathcal{P}^{i}(d_{1,2})\in R_{2}$. Similarly, since $u_{2}d_{2,2}-\xi_{1}^{q}\xi_{3}\in R_{2}$, we have $\mathcal{P}^{i}(u_{2}d_{2,2})\in R_{2}$, proving $\mathcal{P}^{i}(d_{2,2})\in R_{2}$, completing the proof of (a).
The proof of (b) is by induction on $i$. Part (a) gives (b) for $i<q$. Suppose $i=\ell q^{2}+r$ with $r<q^{2}$ and $\ell<q$. Using Lemma 3.5, for $f\in R_{2}$ we have $\mathcal{P}^{i}(f)\in R_{3}$ with $\xi_{3}$-degree at most $\ell$. Since $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$, we have $\mathcal{P}^{i}(u_{2}d_{1,2})\in R_{3}$ with $\xi_{3}$-degree at most $1$ when $\ell=0$ and at most $\ell$ otherwise. Similarly, using the observation $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$, we have $\mathcal{P}^{i}(u_{2}d_{2,2})\in R_{3}$ with $\xi_{3}$-degree at most $1$ when $\ell=0$ and at most $\ell$ otherwise. In the following, let $d$ denote either $d_{1,2}$ or $d_{2,2}$. Using the Cartan identity,
$u_{2}\mathcal{P}^{i}(d)=\mathcal{P}^{i}(u_{2}d)-\sum_{j>0}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d).$
We have seen that $\mathcal{P}^{i}(u_{2}d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell+1$. To prove (b), it is sufficient to show that for $j>0$, we have $u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell+1$. Write $i-j=\ell^{\prime}q^{2}+r^{\prime}$ with $r^{\prime}<q^{2}$ and $\ell^{\prime}\leq\ell$. By induction $u_{2}^{\ell^{\prime}+1}\mathcal{P}^{i}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell^{\prime}+1$.
Suppose $j<q^{2}$. Using Lemma 6.5, we have $\mathcal{P}^{j}(u_{2})/u_{2}\in R_{2}$. Therefore
$u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)=(\mathcal{P}^{i}(u_{2})/u_{2})u_{2}^{\ell+1}\mathcal{P}^{i-j}(d)\in R_{3}$
with $\xi_{3}$-degree at most $\ell+1$, as required.
For $q^{2}\leq j<q^{3}$, we have $\ell^{\prime}+1\leq\ell$ and $\mathcal{P}^{j}(u_{2})\in R_{3}$ with $\xi_{3}$-degree at most $1$. Therefore $u_{2}^{\ell}\mathcal{P}^{i-j}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell$ and $u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell+1$, as required. ∎
Lemma 6.7.
Suppose $i=(q-1)q^{n-2}+r$ with $r<q^{n-2}$ and $i\not\equiv_{(q)}0$. Then $u_{2}^{q}\mathcal{P}^{i}(d_{j,2})$ has $\xi_{3}$-degree at most $q-1$
Proof.
From Lemma 6.6, we know that $u_{2}^{q}\mathcal{P}^{i}(d_{j,2})\in R_{3}$ with with $\xi_{3}$-degree at most $q$. We will show that the $\xi_{3}$-degree at most $q-1$. Using the Cartan identity
$u_{2}^{q}\mathcal{P}^{i}(d_{j,2})=\mathcal{P}^{i}(u_{2}^{q}d_{j,2})-\sum_{k>0}\mathcal{P}^{k}(u_{2})^{q}\mathcal{P}^{i-qk}(d_{j,2}).$
Since $u_{2}^{q}d_{j,2}\in R_{3}$ with $\xi_{3}$-degree $1$, we see that $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})\in R_{3}$ with with $\xi_{3}$-degree at most $q$. However, by Lemma 6.4, the coefficient of $\xi_{3}$ in $u_{2}^{q}d_{j,2}$ is $u_{2}^{q-1}(u_{1}d_{j-1,1})^{q}$, which means that the coefficient of $\xi_{3}\xi_{2}^{q-1}$ in $u_{2}^{q}d_{j,2}$ is $(u_{1}^{q}d_{j-1,1})^{q}$. Therefore the coefficient of $\xi_{3}^{q}$ in $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})$ is $\mathcal{P}^{r}((u_{1}^{q}d_{j-1,1})^{q})$. Since $r\not\equiv_{(q)}0$, $\mathcal{P}^{r}((u_{1}^{q}d_{j-1,1})^{q})=0$ and the $\xi_{3}$-degree of $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})$ is at most $q-1$. Since $i<q^{3}$ and $qk<i$, we have $k<q^{2}$. Thus $\mathcal{P}^{k}(u_{2})/u_{2}\in R_{2}$ and
$\mathcal{P}^{k}(u_{2})^{q}\mathcal{P}^{i-qk}(d_{j,2})=(\mathcal{P}^{k}(u_{2})/u_{2})^{q}u_{2}^{q}\mathcal{P}^{i-qk}(d_{j,2}).$
Write $r=r_{0}+qr^{\prime}$ with $0<r_{0}<q$. If $r^{\prime}=0$, then $i-kp<(q-1)q^{2}$ and $u_{2}^{q}\mathcal{P}^{i-qk}(d_{j,2})$ has $\xi_{3}$-degree at most $q-1$. If $r^{\prime}>0$, then the result follows by induction on $r^{\prime}$. ∎
This completes the proof of part (a) of Theorem 3.11 for $m=2$. Part (b) of Theorem 4.6 for $m=2$ follows from Lemmas 6.4 and 6.1. Part (c) of Theorem 4.6 is just Theorem 3.11.
The following gives part (d) of Theorem 4.6 for $m=2$.
Lemma 6.8.
In $S_{2}^{{{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}}$, $\xi_{3}=(\xi_{2}+c_{2,2})d_{1,2}-\xi_{1}d_{2,2}-c_{3,2}$ with $c_{3,2}\in R_{2}$ and
$-c_{3,2}\equiv_{\langle\xi_{0}^{q+2}\rangle}\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}.$
Proof.
Using Lemma 6.1, $u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}\in R_{1}$. Since $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}$ and $u_{2}d_{2,2}-\xi_{1}^{q}\xi_{3}$ are elements of $R_{2}$,
$u_{2}\xi_{3}-(\xi_{2}+c_{2,2})u_{2}d_{1,2}+\xi_{1}u_{2}d_{2,2}\in R_{2}.$
Since this expression is divisible by $u_{2}$ in $S_{2}$ and lies in $R_{2}$, it is divisible by $u_{2}$ in $R_{2}$. Defining $-c_{3,2}:=(u_{2}\xi_{3}-(\xi_{2}+c_{2,2})u_{2}d_{1,2}-\xi_{1}u_{2}d_{2,2})/u_{2}$ gives $\xi_{3}=(\xi_{2}+c_{2,2})d_{1,2}-\xi_{1}d_{2,2}-c_{3,2}$ with $c_{3,2}\in R_{2}$. Using Lemma 6.4
$(\xi_{2}+c_{2,2})u_{2}d_{1,2}-\xi_{1}u_{2}d_{2,2}$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{3}((\xi_{2}+c_{2,2})\xi_{0}^{q}-\xi_{1}^{q+1})-\xi_{2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$+\xi_{1}(\xi_{2}^{q+1}+\xi_{0}^{q}\xi_{1}^{q^{2}-q}\xi_{2})-c_{2,2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$=$ $u_{2}\xi_{3}+\xi_{0}\xi_{1}^{q}\xi_{2}^{q}+\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}$
$-c_{2,2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
Therefore, using Lemma 6.1 to substitute for $c_{2,2}$,
$-u_{2}c_{3,2}$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $-\xi_{0}\xi_{1}^{q}\xi_{2}^{q}-\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}+c_{2,2}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$\equiv_{\langle\xi_{0}^{q+2}\rangle}$ $-\xi_{0}\xi_{1}^{q}\xi_{2}^{q}-\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}+\xi_{0}\xi_{1}^{q-1}(\xi_{1}\xi_{2}^{q}-\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1})$
$=$ $-\xi_{0}^{2}\xi_{1}^{2q-1}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-q+1}\xi_{2}$
$=$ $-\xi_{1}^{q+1}(\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}).$
Thus $-c_{3,2}\equiv_{\langle\xi_{0}^{q+2}\rangle}\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}.$
Applying $\mathcal{P}^{1}$ to $\xi_{3}=(\xi_{2}+c_{2,2})d_{1,2}-\xi_{1}d_{2,2}-c_{3,2}$ and using $\mathcal{P}^{1}(c_{2,2})=\xi_{1}^{q}$ from Lemma 6.3 gives
$\xi_{2}^{q}=2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}-\mathcal{P}^{1}(c_{3,2})+(\xi_{2}+c_{2,2})\mathcal{P}^{1}(d_{1,2})-\xi_{1}\mathcal{P}^{1}(d_{2,2}).$
Since $c_{3,2}\in R_{2}$, $\mathcal{P}^{1}(c_{3,2})\in R_{2}$. Furthermore, using Lemma 6.8,
$\mathcal{P}^{1}(c_{3,2})\equiv_{\langle\xi_{0}^{q+2}\rangle}\mathcal{P}^{1}(\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2})=2\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}+\xi_{0}^{q}\xi_{1}^{q^{2}-q}$
Using Lemmas 6.4 and 6.5, $\mathcal{P}^{1}(u_{2}d_{1,2})=u_{2}\mathcal{P}^{1}(d_{1,2})$, $\mathcal{P}^{1}(d_{1,2})\in R_{2}$ and
$u_{2}\mathcal{P}^{1}(d_{1,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{0}^{q}\xi_{2}^{q}-2\xi_{0}^{q}\xi_{2}^{q}+\xi_{1}^{q+1}\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{2q}\xi_{2}^{q-2}=\xi_{1}^{q+1}\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{2q}\xi_{2}^{q-2}-\xi_{0}^{q}\xi_{2}^{q}.$
Again using Lemmas 6.4 and 6.5, $\mathcal{P}^{1}(u_{2}d_{2,2})=u_{2}\mathcal{P}^{1}(d_{1,2})$, $\mathcal{P}^{1}(d_{2,2})\in R_{2}$ and $u_{2}\mathcal{P}^{1}(d_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{2}^{q}-\xi_{1}\xi_{2}^{q}-\xi_{0}^{q}\xi_{1}^{q^{2}}=-\xi_{0}^{q}\xi_{1}^{q^{2}}$. Since $u_{2}\equiv_{\langle\xi_{0}^{q+1}\rangle}-\xi_{1}^{q+1}+\xi_{0}^{q}\xi_{2}$, we have
$\mathcal{P}^{1}(d_{1,2})\equiv_{\langle\xi_{0}^{q+1}\rangle}-\xi_{2}^{q-1}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2}$
and $\mathcal{P}^{1}(d_{2,2})\equiv_{\langle\xi_{0}^{q+1}\rangle}\xi_{0}^{q}\xi_{1}^{q^{2}-q-1}$. Since $\mathcal{P}^{1}(d_{1,2})$ and $\mathcal{P}^{1}(d_{2,2})$ lie in $R_{2}$, we have $\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}\in R_{2}$. Using the congruences we have
$\xi_{2}^{q}$ $\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}+(2\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}-\xi_{0}^{q}\xi_{1}^{q^{2}-q})$
$-(\xi_{2}+c_{2,2})(\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2})+\xi_{0}^{q}\xi_{1}^{q^{2}-q}$
$\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}+3\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{2}^{q}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}$
$-c_{2,2}(\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2})$
$\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}-\xi_{2}^{q}+3\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{2}\xi_{1}^{2q-2}\xi_{2}^{q-2}$
$-\xi_{0}\xi_{1}^{q-1}(\xi_{2}^{q-1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-2})$
$\equiv_{\langle\xi_{0}^{q+1}\rangle}$ $2\xi_{1}^{q}d_{1,2}-2\xi_{0}^{q}d_{2,2}-\xi_{2}^{q}+2\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}.$
Therefore
$\xi_{2}^{q}\equiv_{\langle\xi_{0}^{q+1}\rangle}\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}.$
Recall that $-\xi_{1}^{q}d_{1,2}+\xi_{0}^{q}d_{2,2}\in R_{2}$ and define $r_{2,1}\in R_{2}$ by
(3) $\xi_{0}^{q+1}r_{2,1}=\xi_{2}^{q}-\xi_{1}^{q}d_{1,2}+\xi_{0}^{q}d_{2,2}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}.$
Then
(4) $\xi_{2}^{q}=\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}-\gamma_{2,2}^{(1)}$
with $\gamma_{2,2}^{(1)}=-\xi_{0}^{q+1}r_{2,1}-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}\in R_{2}$ and $\nu(\gamma_{2,2}^{(1)})>q+1$. Thus part (e) of Theorem 4.6 is satisfied for $m=2$.
Lemma 7.2 does not rely on induction and the proof is valid for $m=2$. This completes the proof of Theorem 4.6 for $m=2$.
Remark 6.9.
Notice that in the process of deriving Equation (4), the expression $-\xi_{2}^{q}$ arises on the right hand side. If the coefficent had been +1 rather than -1 then this process would not have provided a formula for $\xi_{2}^{q}$. It is for this reason that naively applying a Steenrod operation to the formula for $\xi_{n-1-i}^{q^{i}}$ is not guaranteed to give a formula for $\xi_{n-2-i}^{q^{i+1}}$.
7. Proof of the Main Theorem
The proof is by induction on $m$. The base case, $m=2$, is given in Section 6. For the induction step, we will assume the theorem is true for $S_{\ell}^{G_{\ell}}$ for $2\leq\ell<m$ and, in the proof of a given part of the theorem, we will assume the earlier parts are true for $S_{m}^{G_{m}}$.
Remark 7.1.
Part (c) is Theorem 3.11.
Lemma 7.2.
Part (f) of Theorem 4.6 follows from Parts (a)–(e).
Proof.
Let $A$ denote the subalgebra of $S_{m}^{G_{m}}$ generated by $\mathcal{H}\cup\{\xi_{m},\ldots,\xi_{n-2}\}$. Using Part (c), $A\subseteq S_{m}^{G_{m}}$ is an integral extension. Using Part (d), $R_{n-1}\subset A$. Therefore $\mathcal{F}(A)=\mathcal{F}(R_{n-1})=\mathcal{F}(S_{m}^{G_{m}})$. Thus to show $A=S_{m}^{G_{m}}$, it is sufficient to show that $A$ is integrally closed in its field of fractions. We will use [17, Proposition 1.1] to prove that $A$ is a UFD. Recall that a UFD is integrally closed in its field of fractions (see, for example, [13, Proposition 4.10]).
Using the definition and part (e), $A$ is the $\mathbb{F}_{q}[\mathcal{H}]$-module generated by $\mathcal{B}$. Note that the number of elements of the set $\mathcal{B}$ is $\prod_{i=1}^{m-1}q^{i}=q^{m(m-1)/2}$. We know that $S_{m}^{G_{m}}$ is a free $\mathbb{F}_{q}[\mathcal{H}]$-module of rank
$\frac{\prod_{i=0}^{m-1}(q^{i}+1)\prod_{j=1}^{m}(q^{n-1}-q^{n-1-j})}{2q^{m(m-1)}(q^{m}-1)\prod_{k=1}^{m-1}(q^{2k}-1)},$
which simplifies to $q^{m(m-1)/2}$. Therefore the degree of the field extension $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])\subset\mathcal{F}(S_{m}^{G_{m}})$ is $q^{m(m-1)/2}$. Since $\mathcal{F}(A)=\mathcal{F}(S_{m}^{G_{m}})$, the degree of the field extension $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])\subset\mathcal{F}(A)$ is $q^{m(m-1)/2}$. We know that $\mathcal{B}$ is a spanning set for $\mathcal{F}(A)$ over $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])$. By comparing the order of $\mathcal{B}$ with the degree of the extension, we see that $\mathcal{B}$ is linearly independent over $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])$. Therefore $A$ is a free $\mathbb{F}_{q}[\mathcal{H}]$-module with basis $\mathcal{B}$ and $A$ is Cohen-Macaulay. By comparing Hilbert series, we see that $A$ is a complete intersection with relations given by Part (e). Therefore, to complete the proof of Part (f), we need only show that $A=S_{m}^{G_{m}}$.
From Part (b) for $m$ and $m-1$, we have $\operatorname{LT}(u_{m})=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1}+q^{m-2}+\cdots+1}$ and $\operatorname{LT}(u_{m-1})=(-1)^{\lfloor(m-1)/2\rfloor}\xi_{m-2}^{q^{m-2}+q^{m-3}+\cdots+1}$. Therefore, since $\mathcal{H}$ is an hsop (see Part (c)), $\{u_{m-1},u_{m}\}$ is a partial hsop. Since $A$ is Cohen-Macaulay, this means that $[u_{m-1},u_{m}]$ is a regular sequence in $A$. From (a) for $m-1$, we have $\ker(\Phi_{n-4,m-2})=u_{m-1}R_{n-4}$. Therefore $u_{m-1}$ is prime in $R_{n-4}$. Since $R_{n-1}$ is a polynomial extension of $R_{n-4}$, $u_{m-1}$ is prime in $R_{n-1}$ and in its localisation $R_{n-1}[u_{m}^{-1}]$. Using Part (d), $R_{n-1}\subset A$ and by definition $u_{m}d_{i,m}\in R_{n-1}$. Therefore $A[u_{m}^{-1}]=R_{n-1}[u_{m}^{-1}]$ and $u_{m-1}$ is prime in $A[u_{m}^{-1}]$. To satisfy the hypotheses of [17, Proposition 1.1] and complete the proof, we need to show that $A[u_{m-1}^{-1}]$ is a UFD.
From Part (a), we have $u_{m}d_{1,m}-u_{m-1}^{q}\xi_{n-1}\in R_{n-2}$. Therefore $R_{n-1}\subset R_{n-2}[d_{1,m},u_{m-1}^{-1}]$ and $\mathcal{F}(R_{n-1})=\mathcal{F}(R_{n-2}[d_{1,m}])$. Hence $R_{n-2}[d_{1,m}]$ is a polynomial algebra and $R_{n-2}[d_{1,m},u_{m-1}^{-1}]$ is a UFD. We prove $A[u_{m-1}^{-1}]$ is a UFD by showing $A[u_{m-1}^{-1}]=R_{n-2}[d_{1,m},u_{m-1}^{-1}]$. By definition, $R_{n-2}[d_{1,m}]\subset A$. From Part (a), we have $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$ for $i>1$. Cross-multiplying to eliminate $\xi_{n-1}$ gives
$u_{m-1}^{q}u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}u_{m}d_{1,m}\in R_{n-2}.$
Using $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$ (from(a)), we see that
$u_{m-1}^{q}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}d_{1,m}\in R_{n-2}$
(with $u_{m-1}d_{i-1,m-1}\in R_{n-3}$). Thus $d_{i,m}\in R_{n-2}[d_{1,m},u_{m-1}^{-1}]$, as required. ∎
Lemma 7.3.
Part (d) of Theorem 4.6 follows from Part (a).
Proof.
From (a) we have
$u_{m}=(\xi_{n-2}+c_{n-2,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(u_{m-1}d_{i,m-1})^{q}$
with $c_{j,m}\in R_{j-1}$ and $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$. Therefore
$\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-1-i}+c_{n-1-i,m})(u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1})\in R_{n-2}.$
Note that the coefficient of $\xi_{n-1}$ in this expression is $u_{m}$. Dividing by $u_{m}$ gives
$c_{n-1,m}:=\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-1-i}+c_{n-1-i,m})d_{i,m}\right)-\xi_{n-1}\in R_{n-2}$
as required. ∎
7.4. Proof of Part (a)
The proof is by induction on $m$. The base cases $m=2$ is given in Section 6. By the induction hypothesis $S_{m-1}^{G_{m-1}}$ is generated by $\{\xi_{0},\ldots,\xi_{n-4},d_{1,m-1},\ldots,d_{m-1,m-1}\}$. The Hook subgroup for $G_{m}$ is a normal subgroup of the pointwise stabiliser $(G_{m})_{x_{1}}$. If we identify $G_{m-1}$ with the pointwise stabiliser of $\{x_{1},y_{1}\}$ in $G_{m}$, then $(G_{m})_{x_{1}}=G_{m-1}H$. Arguing as in the proof of Theorem 5.9 and using Theorem 5.4, we see that $S_{m}^{(G_{m})_{x_{1}}}$ is generated by
$\{N(y_{1}),x_{1},\xi_{0},\ldots,\xi_{n-3},\psi_{1}(\sigma(d_{1,m-1})),\ldots\psi_{1}(\sigma(d_{m-1,m-1}))\}.$
In this context, the compliance relation (Lemma 5.2) allows us to rewrite $\xi_{n-2}$ in terms of the generators of $S_{m}^{(G_{m})_{x_{1}}}$. For degree reasons, the resulting expression is linear in the $\psi_{1}(\sigma(d_{j,m-1}))$, giving
$\xi_{n-2}=x_{1}N(y_{1})+\lambda_{0}+\sum_{j=1}^{m-1}\lambda_{j}\psi_{1}(\sigma(d_{j,m-1}))$
with $\lambda_{j}\in R_{n-3}[x_{1}]$. Since $u_{m}=x_{1}N(y_{1})\psi_{1}(\sigma(u_{m-1}))$, we have
$u_{m}=\psi_{1}(\sigma(u_{m-1}))\left(\xi_{n-2}-\lambda_{0}-\sum_{j=1}^{m-1}\lambda_{j}\psi_{1}(\sigma(d_{j,m-1}))\right).$
We know that $\psi_{1}$ and $\sigma$ are algebra maps. Furthermore, $\psi_{1}(\sigma(f))=\psi_{1}(f)$ for $f\in R_{n-3}$. By induction $u_{m-1}$ and $u_{m-1}d_{j,m-1}$lie in $R_{n-3}$. Therefore
$u_{m}=(\xi_{n-2}-\lambda_{0})\psi_{1}(u_{m-1})-\sum_{j=1}^{m-1}\lambda_{j}\psi_{1}(u_{m-1}d_{j,m-1}).$
Write $\lambda_{j}=\overline{\lambda_{j}}+x_{1}\lambda_{j}^{\prime}$ with $\overline{\lambda_{j}}\in R_{n-3}$ and $\lambda_{j}^{\prime}\in R_{n-3}[x_{1}]$ and observe that $\psi_{1}(\lambda_{j})=\psi_{1}(\overline{\lambda_{j}})$. Using Lemma 4.2, $\psi_{1}(\psi_{1}(f))=\psi_{1}(f)^{q}$ for $f\in R_{n-3}$. Applying $\psi_{1}$ to the above expression and using the fact that $\psi_{1}(u_{m})=0$ gives
$0=\psi_{1}(\xi_{n-2}-\overline{\lambda_{0}})\psi_{1}(u_{m-1})^{q}-\sum_{j=1}^{m-1}\psi_{1}(\overline{\lambda_{j}})\psi_{1}(u_{m-1}d_{j,m-1})^{q}.$
Therefore
$(\xi_{n-2}-\overline{\lambda_{0}})u_{m-1}^{q}-\sum_{j=1}^{m-1}\overline{\lambda_{j}}(u_{m-1}d_{j,m-1})^{q}\in R_{n-2}\cap\ker(\psi_{1}).$
Using Lemma 4.2, $u_{m}$ divides every element of $R_{n-2}\cap\ker(\psi_{1})$. By comparing lead terms, we conclude
$u_{m}=(\xi_{n-2}-\overline{\lambda_{0}})u_{m-1}^{q}-\sum_{j=1}^{m-1}\overline{\lambda_{j}}(u_{m-1}d_{j,m-1})^{q}\in R_{n-2}$
and $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$. Applying $\mathcal{P}^{e(i,m)}$ to this expression shows that $u_{m}d_{i,m}=\mathcal{P}^{e(i,m)}(u_{m})\in R_{n-1}$ and $u_{m}d_{i,m}-\xi_{n-1}\mathcal{P}^{e(i,m)-q^{n-2}}(u_{m-1}^{q})\in R_{n-2}$. Since $e(i,m)-q^{n-2}=qe(i-1,m-1)$, we have $u_{m}d_{i,m}-\xi_{n-1}(u_{m-1}d_{i-1,m-1})^{q}\in R_{n-2}$.
As an element of $S_{m}$, $u_{m}$ is a product of linear forms. Therefore $u_{m}$ divides $\mathcal{P}^{i}(u_{m})$ in $S_{m}$. If $\mathcal{P}^{i}(u_{m})\in R_{n-2}$, then $\mathcal{P}^{i}(u_{m})\in\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$ and $u_{m}$ divides $\mathcal{P}^{i}(u_{m})$ in $R_{n-2}$. For $i<q^{n-2}$, $\mathcal{P}^{i}(u_{m})\in R_{n-2}$ and, therefore, $\mathcal{P}^{i}(u_{m})/u_{m}\in R_{n-2}$ By induction, $\mathcal{P}^{i}(u_{m-1})=0$ for $0<i<q^{m-2}$. Therefore, for $0<i<q^{m-1}$, $\mathcal{P}^{i}(u_{m-1}^{q})=0$ and $\mathcal{P}^{i}(u_{m})\in R_{n-3}$. Since $u_{m}$ is linear in $\xi_{n-2}$ and divides $\mathcal{P}^{i}(u_{m})\in R_{n-3}$, we conclude that $\mathcal{P}^{i}(u_{m})=0$ for $0<i<q^{m-1}$.
For $i<q^{n-1}$, write $i=\ell q^{n-2}+r$ with $\ell<q$ and $r<q^{n-2}$. We will prove by induction on $i$ that $u_{m}^{\ell+1}\mathcal{P}^{i}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$. For $i=0$, we have $\ell=0$ and $u_{m}\mathcal{P}^{0}(d_{j,m})=u_{m}d_{j,m}\in R_{n-1}$ with $\xi_{n-1}$-degree $1$. For $0<i<q^{m-1}$, we have $u_{m}\mathcal{P}^{i}(d_{j,m})=\mathcal{P}^{i}(u_{m}d_{j,m})\in R_{n-2}$, giving $\mathcal{P}^{i}(d_{j,m})\in R_{n-2}$. For $i\geq q^{m-1}$, using the Cartan identity gives
$u_{m}\mathcal{P}^{i}(d_{j,m})=\mathcal{P}^{i}(u_{m}d_{j,m})-\sum_{k>0}\mathcal{P}^{k}(u_{m})\mathcal{P}^{i-k}(d_{j,m}).$
Since $u_{m}d_{j,m}\in R_{n-1}$ with $\xi_{n-1}$-degree $1$, we see that $\mathcal{P}^{i}(u_{m}d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$. For $k>0$, write $i-k=\ell^{\prime}q^{n-2}+r^{\prime}$ with $\ell^{\prime}\leq\ell$ and $r^{\prime}<q^{n-2}$. Since $i-k<i$, the induction hypothesis gives $u_{m}^{\ell^{\prime}+1}\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell^{\prime}+1$. If $k<q^{n-2}$, then $\mathcal{P}^{k}(u_{m})/u_{m}\in R_{n-2}$ and
$u_{m}^{\ell}\mathcal{P}^{k}(u_{m})\mathcal{P}^{i-k}(d_{j,m})=(\mathcal{P}^{k}(u_{m})/u_{m})u_{m}^{\ell+1}\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$
with $\xi_{n-1}$-degree at most $\ell+1$, as required. Suppose $q^{n-2}\leq k$. Then $\ell^{\prime}<\ell$ and $u_{m}^{\ell}\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell$. Since $\mathcal{P}^{k}(u_{m})\in R_{n-1}$ with with $\xi_{n-1}$-degree at most $1$, $u_{m}^{\ell}\mathcal{P}^{k}(u_{m})\mathcal{P}^{i-k}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$, as required.
Suppose $i=(q-1)q^{n-2}+r$ with $r<q^{n-2}$ and $i\not\equiv_{(q)}0$. We know that $u_{m}^{q}\mathcal{P}^{i}(d_{j,m})\in R_{n-1}$ with with $\xi_{n-1}$-degree at most $q$. We will show that the $\xi_{n-1}$-degree at most $q-1$. Using the Cartan identity
$u_{m}^{q}\mathcal{P}^{i}(d_{j,m})=\mathcal{P}^{i}(u_{m}^{q}d_{j,m})-\sum_{k>0}\mathcal{P}^{k}(u_{m})^{q}\mathcal{P}^{i-qk}(d_{j,m}).$
Since $u_{m}^{q}d_{j,m}\in R_{n-1}$ with $\xi_{n-1}$-degree $1$, we see that $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})\in R_{n-1}$ with with $\xi_{n-1}$-degree at most $q$. However, the coefficient of $\xi_{n-1}$ in $u_{m}^{q}d_{j,m}$ is given by $u_{m}^{q-1}(u_{m-1}d_{j-1,m-1})^{q}$, which means that the coefficient of $\xi_{n-1}\xi_{n-2}^{q-1}$ in $u_{m}^{q}d_{j,m}$ is $(u_{m-1}^{q}d_{j-1,m-1})^{q}$. Therefore the coefficient of $\xi_{n-1}^{q}$ in $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})$ is given by $\mathcal{P}^{r}((u_{m-1}^{q}d_{j-1,m-1})^{q})$. Since $r\not\equiv_{(q)}0$, $\mathcal{P}^{r}((u_{m-1}^{q}d_{j-1,m-1})^{q})=0$ and the $\xi_{n-1}$-degree of $\mathcal{P}^{i}(u_{m}^{q}d_{j,m})$ is at most $q-1$. Since $i<q^{n-1}$ and $qk<i$, we have $k<q^{n-2}$. Thus $\mathcal{P}^{k}(u_{m})/u_{m}\in R_{n-2}$ and
$\mathcal{P}^{k}(u_{m})^{q}\mathcal{P}^{i-qk}(d_{j,m})=(\mathcal{P}^{k}(u_{m})/u_{m})^{q}u_{m}^{q}\mathcal{P}^{i-qk}(d_{j,m}).$
Write $r=r_{0}+qr^{\prime}$ with $0<r_{0}<q$. If $r^{\prime}=0$, then $i-kp<(q-1)q^{n-2}$ and $u_{m}^{q}\mathcal{P}^{i-qk}(d_{j,m})$ has $\xi_{n-1}$-degree at most $q-1$. If $r^{\prime}>0$, then the result follows by induction on $r^{\prime}$.
Using the induction hypothesis, in $S_{m-1}^{G_{m-1}}$, we have
$\xi_{n-3}=\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-3-i}+c_{n-3-i,m-1})d_{i,m-1}\right)-c_{n-3,m-1}.$
Using Lemma 3.6,
$\mathcal{P}^{q^{n-3}}(\xi_{n-3-i}d_{i,m-1})=\xi_{n-2-i}d_{i,m-1}^{q}+\xi_{n-3-i}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}$
and
$\mathcal{P}^{q^{n-3}}(c_{n-3-i,m-1}d_{i,m-1})=\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1})d_{i,m-1}^{q}+c_{n-3-i,m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}$
where $a_{i}=\deg(d_{i,m-1})-1$. Therefore, applying $\mathcal{P}^{q^{n-3}}$ to the relation above gives
$\xi_{n-2}$ $=$ $\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1}))d_{i,m-1}^{q}\right)$
$+\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\right)-\mathcal{P}^{q^{n-3}}(c_{n-3,m-1}).$
Multiplying by $u_{m-1}^{q}$ gives
$u_{m-1}^{q}\xi_{n-2}$ $=$ $\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1}))(u_{m-1}d_{i,m-1})^{q}\right)$
$+\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}u_{m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\right)$
$-u_{m-1}^{q}\mathcal{P}^{q^{n-3}}(c_{n-3,m-1}).$
By induction $u_{m-1}d_{i,m-1}\in R_{n-3}$ with $\xi_{n-3}$-degree $1$ and, since $a_{i}=q^{n-3}-q^{n-3-i}-1$, $u_{m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\in R_{n-3}$ with $\xi_{n-3}$-degree at most $q-1$. Therefore the right hand side of the relation is an element of $R_{n-3}$ and the relation lifts to give an element of $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}\subset S_{m}^{G_{m}}$. Comparing lead terms gives
$u_{m}$ $=$ $u_{m-1}^{q}\xi_{n-2}-\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-2-i}+\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1}))(u_{m-1}d_{i,m-1})^{q}\right)$
$-\left(\sum_{i=1}^{m-1}(-1)^{i+1}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}u_{m-1}^{q}\mathcal{P}^{a_{i}}d_{i,m-1}\right)$
$+u_{m-1}^{q}\mathcal{P}^{q^{n-3}}(c_{n-3,m-1}).$
Comparing the two expressions for $u_{m}$ and using the fact that $u_{m-1}^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})$ has $\xi_{n-3}$-degree at most $q-1$ while $(u_{m-1}d_{i,m-1})^{q}$ has $\xi_{n-3}$-degree $q$, we conclude that
$u_{m}=(\xi_{n-2}+c_{n-2,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(u_{m-1}d_{i,m-1})^{q}$
with
$c_{n-2,m}=\mathcal{P}^{q^{n-3}}(c_{n-3,m-1})+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-3-i}+c_{n-3-i,m-1})^{q}\mathcal{P}^{a_{i}}(d_{i,m-1})\in R_{n-3}$
and $c_{n-2-i,m}=\mathcal{P}^{q^{n-3-i}}(c_{n-3-i,m-1})\in R_{n-3-i}$ for $0<i<m$. This completes the proof of part (a). ∎
7.5. Proof of Part (b)
The proof is by induction on $m$. For $m=2$, the result follows from Lemmas 6.1 and 6.4.
Suppose $m>2$. By induction, $u_{m-1}=M(0,m-1)+\delta_{0,m-1}$ and $u_{m-1}d_{i,m-1}=M(i,m-1)+\delta_{i,m-1}$. Substituting into the expression for $u_{m}$ from part (a) gives
$u_{m}$ $=$ $(\xi_{n-2}+c_{n-2,m})(M(0,m-1)+\delta_{0,m-1})^{q}$
$+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(M(i,m-1)+\delta_{i,m-1})^{q}$
$=$ $\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}M(i,m-1)^{q}+\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}\delta_{i,m-1}^{q}$
$+\sum_{i=0}^{m-1}(-1)^{i}c_{n-2-i}(M(i,m-1)^{q}+\delta_{i,m-1}^{q}).$
Observe that $M(0,m)=\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}M(i,m-1)^{q}$ and define
$\delta_{0,m}=\sum_{i=0}^{m-1}(-1)^{i}\xi_{n-2-i}\delta_{i,m-1}^{q}+\sum_{i=0}^{m-1}(-1)^{i}c_{n-2-i}(M(i,m-1)^{q}+\delta_{i,m-1}^{q}).$
Then $u_{m}=M(0,m)+\delta_{0,m}$ with $\delta_{0,m}\in R_{n-2}$. Since $\nu(c_{n-2-i})>1$, $\nu(M(i,m-1))=1+q+\cdots+q^{m-2}$, and $\nu(\delta_{i,m-1})>1+q+\cdots+q^{m-2}$, we have $\nu(\delta_{0,m})>1+q+\cdots+q^{m-1}$.
Apply $\mathcal{P}^{e(i,m)}$ to $u_{m}$ to get $u_{m}d_{i,m}=\mathcal{P}^{e(i,m)}(M(0,m))+\mathcal{P}^{e(i,m)}(\delta_{0,m})$. Define $\delta_{i,m}:=u_{m}d_{i,m}-M(i,m)$ so that $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$, and observe that $\delta_{i,m}\in R_{n-1}$. Using Lemma 4.3,
$\nu(\mathcal{P}^{e(i,m)}(\delta_{0,m}))\geq\nu(\delta_{0,m})>1+q+\cdots+q^{m-1}.$
Using Lemma 4.5, $\nu(\mathcal{P}^{e(i,m)}(M(0,m)-M(i,m))>1+q+\cdots+q^{m-1}$. Therefore $\nu(\delta_{i,m})>1+q+\cdots+q^{m-1}$.
It is clear that using the weighted grevlex order on $R_{n-1}$,
$\operatorname{LT}(M(0,m))=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1}+q^{m-2}+\cdots+1}$
and
$\operatorname{LT}(M(i,m))=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1-i}+\cdots+1}\xi_{m}^{q^{m-1}+\cdots+q^{m-i}}.$
From Lemma 4.4, $\operatorname{LT}(\delta_{i,m})<\operatorname{LT}(M(i,m))$ for $0\leq i\leq m$. Therefore $\operatorname{LT}(u_{m})=\operatorname{LT}(M(0,m))$ and $\operatorname{LT}(u_{m}d_{im})=\operatorname{LT}(M(i,m))$, completing the proof part (b). ∎
7.6. Proof of Part (e)
The proof is by induction on $m$. For $m=2$, Equation 3 gives
$\xi_{2}^{q}=\xi_{1}^{q}d_{1,2}-\xi_{0}^{q}d_{2,2}+\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}+\xi_{0}^{q+1}r_{2,1}$
with $r_{2,1}\in R_{2}$. Define $\gamma_{1,2}^{(1)}=\gamma_{0,2}^{(1)}=0$ and $\gamma_{2,2}^{(1)}=-\xi_{0}\xi_{1}^{q-1}\xi_{2}^{q-1}-\xi_{0}^{q+1}r_{2,1}$. Since $\nu(\gamma_{2,2}^{(1)})>q+1$, the result holds for $m=2$.
Suppose $m>2$. By induction we have parts (d) and (e) of the theorem for $S_{m-1}^{G_{m-1}}$. Therefore, for $1\leq i\leq m$, we have the relation
$0=u_{m-1}\left(\xi_{n-2-i}^{q^{i-1}}+\gamma_{n-2-i,m-1}^{(i-1)}-\sum_{j=1}^{m-1}(-1)^{j+1}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})d_{j,m-1}\right)$
in $R_{n-3}$. Define
$\gamma:=\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}u_{m}d_{j+1,m}.$
By part (a), $u_{m}d_{j,m}-(u_{m-1}d_{j-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$ for $1\leq j\leq m$, using the convention $d_{0,m-1}=1$. Therefore $\gamma\in R_{n-1}$. Furthermore $\gamma$ is linear in $\xi_{n-1}$ with coefficient
$\left(-\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})u_{m-1}d_{j,m-1}\right)^{q}=0.$
Hence $\gamma\in R_{n-2}$. From part (a), $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$. Since $u_{m}$ divides $\gamma$ in $S_{m}^{G_{m}}$ and $\gamma$ lies in $R_{n-2}$, we see that $\gamma\in u_{m}R_{n-2}$. Thus $\gamma/u_{m}\in R_{n-2}$.
Define an $(m+1)\times(m+1)$ matrix in $R_{n-1}$ by
$M_{m}^{(i)}:=\begin{pmatrix}\xi_{n-1-i}^{q^{i}}&\cdots&\xi_{m-1-i}^{q^{i}}\\ &M_{m}&\end{pmatrix}.$
Since row $1$ and row $i+1$ of $M_{m}^{(i)}$ are equal, $\det(M_{m}^{(i)})=0$. Therefore, computing the determinant by expanding across the first row gives
$\sum_{j=0}^{m}(-1)^{j}\xi_{n-1-i-j}^{q^{i}}M(j,m)=0.$
Hence
$\xi_{n-1-i}^{q^{i}}M(0,m)=\sum_{j=1}^{m}(-1)^{j+1}\xi_{n-1-i-j}^{q^{i}}M(j,m).$
From part (b), $u_{m}d_{j,m}=M(j,m)+\delta_{j,m}$ with $\nu(\delta_{j,m})>1+q+\cdots+q^{m-1}$. Substituting into the expression for $\gamma$ gives
$\gamma$ $=$ $\sum_{j=0}^{m-1}(-1)^{j}\xi_{n-2-i-j}^{q^{i}}M(j+1,m)+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+$ $\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}$
$=$ $\xi_{n-1-i}^{q^{i}}M(0,m)+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+$ $\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}.$
Therefore
$\gamma-\xi_{n-1-i}^{q^{i}}u_{m}$ $=$ $-\xi_{n-1-i}^{q^{i}}\delta_{0,m}+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+$ $\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}.$
By induction, $\nu(\gamma_{j,m-1}^{(i-1)})>q^{i-1}$. Therefore
$\nu(\gamma-\xi_{n-1-i}^{q^{i}}u_{m})>q^{i}+1+q+\cdots+q^{m-1},$
giving $\nu(\gamma/u_{m}-\xi_{n-1-i}^{q^{i}})>q^{i}$. Define $\gamma_{n-1-i,m}^{(i)}:=\gamma/u_{m}-\xi_{n-1-i}^{q^{i}}$ and, for $m-1-i\leq\ell<n-1-i$, define $\gamma_{\ell,m}^{(i)}:=(\gamma_{\ell,m-1}^{(i-1)})^{q}$. Then
$\gamma/u_{m}=\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i}}+\gamma_{n-2-i-j,m}^{(i)})d_{j+1,m}$
and
$\xi_{n-1-i}^{q^{i}}$ $=$ $\gamma/u_{m}-\gamma_{n-1-i,m}^{(i)}$
$=$ $\left(\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i}}+\gamma_{n-2-i-j,m}^{(i)})d_{j+1,m}\right)-\gamma_{n-1-i,m}^{(i)}$
$=$ $\left(\sum_{j=1}^{m}(-1)^{j+1}(\xi_{n-1-i-j}^{q^{i}}+\gamma_{n-1-i-j,m}^{(i)})d_{j,m}\right)-\gamma_{n-1-i,m}^{(i)}$
with $\nu(\gamma_{k,m}^{(i)})>q^{i}$, as required. ∎
8. Generation Over The Steenrod Algebra
Let $\mathcal{A}$ denote the algebra of operations on $S_{m}$ generated by the Steenrod operations with the product given by composition.
Theorem 8.1.
$S_{m}^{G_{m}}$ is generated as an $\mathcal{A}$-algebra by the two elements $\xi_{0}$ and $d_{1,m}$.
Proof.
Let $C$ denote the smallest algebra, closed under the action of the Steenrod algebra and containing $\{\xi_{0},d_{1,m}\}$. Clearly $C\subseteq S_{m}^{G_{m}}$. Since $\mathcal{P}^{q^{i}}(\xi_{i})=\xi_{i+1}$ we see that $\xi_{i}\in C$ for all $i\geq 0$. It remains to show that $d_{i,m}\in C$ for $i=1,2,3,\dots,m$. We proceed by induction. Suppose then that $d_{1,m},d_{2,m},\dots,d_{i-1,m}\in C$ and consider $F:=\mathcal{P}^{q^{m-i-1}}(d_{i-1,m})$. Recall, from the proof of Theorem 3.1, that $d_{i-1,m}\equiv_{I}\pm d_{i-1}(W_{1})$ where $I$ is the ideal $I=\langle x_{1},x_{2},\ldots,x_{m}\rangle$, $W_{1}={\rm Span}_{\mathbb{F}_{q}}\{y_{1},\ldots,y_{m}\}$, and $d_{i-1}(W_{1})$ is the $(i-1)^{\text{th}}$ Dickson invariant for $W_{1}$. Then $F\in S_{m}^{G_{m}}$ is homogeneous of degree $q^{m-1}-q^{m-i}$. Since $\mathcal{P}^{q^{m-i-1}}(d_{i-1}(W_{1}))=-d_{i}(W_{1})$ (see [24]) it follows that $F\equiv_{I}\mathcal{P}^{q^{m-i-1}}(d_{i-1}(W_{1}))=-d_{i}(W_{1})\neq 0$. But the only invariants of degree $q^{m-1}-q^{m-i}$ which do not vanish on $W_{1}$ are scalar multiples of $d_{i,m}$. Hence $F=\pm d_{i,m}+f$ where $f\equiv_{I}0$ and $f\in\mathbb{F}_{q}[\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{i-1,m}]\subseteq C$. This shows that $d_{i,m}\in C$ as required. ∎
Here is an alternate proof of the theorem.
Proof.
As in the proof of Theorem 3.11, let $I$ denote the ideal in $S_{m}$ generated by $\{x_{1},\ldots,x_{m}\}$. Let $J$ denote the ideal in $S_{m}^{G_{m}}$ generated by $\{\xi_{0},\ldots,\xi_{n-2}\}$. Since $J\subset I$, the inclusion of $S_{m}^{G_{m}}$ into $S_{m}$ induces a map from $S_{m}^{G_{m}}/J$ to $S_{m}/I$. Since both $I$ and $J$ are closed under the action of the Steenrod algebra, this is a map of $\mathcal{A}$-algebras. Let $D$ denote the subalgebra of $S_{m}$ generated by $\{d_{1,m},\ldots,d_{m,m}\}$, let $\widetilde{D}$ denote the image of $D$ in $S_{m}^{G_{m}}/J$ and let $\overline{D}$ denote the image of $D$ in $S_{m}/I$. As algebras $D$, $\widetilde{D}$, and $\overline{D}$ are isomorphic. Therefore the map from $S_{m}^{G_{m}}/J$ to $S_{m}/I$ restricts to an isomorphism $\widetilde{D}$ to $\overline{D}$ as $\mathcal{A}$-algebras. From the proof of Theorem 3.9, $\overline{D}$ is isomorphic to $(\mathbb{F}_{q}[W_{1}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}})^{q^{m-1}}$. Since $d_{1}(W_{1})$ generates $\mathbb{F}_{q}[W_{1}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}}$ (see [24]), $\overline{d_{1}}$ generates $\overline{D}$ and $\tilde{d_{1}}$ generates $\widetilde{D}$, as $\mathcal{A}$-algebras. From this we see that $\{\xi_{0},d_{1,m}\}$ generates $S_{m}^{G_{m}}$ as an $\mathcal{A}$-algebra. ∎
Corollary 8.2.
$S_{m}^{G_{m}}$ is generated as an $\mathcal{A}$-algebra by $\{\xi_{0},\mathcal{R}(N(y_{1})^{q-1})\}$.
Proof.
Since $\deg(\mathcal{R}(N(y_{1})^{q-1}))=\deg(d_{1,m})<\deg(d_{i,m})$ for $i>1$, it follows from Theorem 4.6, that $\mathcal{R}(N(y_{1})^{q-1})=cd_{1,m}+\delta$ for some scalar $c$ and $\delta\in R_{n-2}$. By Theorem 5.21 and Lemma 3.10
$\operatorname{LT}(\mathcal{R}(N(y_{1})^{q-1}))=y_{1}^{(q-1)q^{n-2}}=\operatorname{LT}(d_{1,m})$
using the lexicographic order on $S_{m}$. Since $\delta\in I$, we see that $c=1$, and the result follows from Theorem 8.1. ∎
Acknowledgement
The computer algebra package Magma [2] was very useful in this work. It was used to test and confirm many hypotheses in low dimensions for small primes. We are grateful to Karl Dilcher and Keith Taylor (Canadian Mathematical Society book editors) and Donna Chernyk (Springer) for encouraging us to work on this problem.